### v2

# Content-based image retrieval for multi-class volumetric radiology images: a benchmark study

Abstract
While content-based image retrieval (CBIR) has been extensively studied in natural image retrieval, its application to medical images presents ongoing challenges, primarily due to the 3D nature of medical images. Recent studies have shown the potential use of pre-trained vision embeddings for CBIR in the context of radiology image retrieval. However, a benchmark for the retrieval of 3D volumetric medical images is still lacking, hindering the ability to objectively evaluate and compare the efficiency of proposed CBIR approaches in medical imaging. In this study, we extend previous work and establish a benchmark for region-based and localized multi-organ retrieval using the TotalSegmentator dataset (TS) with detailed multi-organ annotations. We benchmark embeddings derived from pre-trained supervised models on medical images against embeddings derived from pre-trained unsupervised models on non-medical images for 29 coarse and 104 detailed anatomical structures in volume and region levels. For volumetric image retrieval, we adopt a late interaction re-ranking method inspired by text matching. We compare it against the original method proposed for volume and region retrieval and achieve a retrieval recall of 1.0 for diverse anatomical regions with a wide size range. The findings and methodologies presented in this paper provide insights and benchmarks for further development and evaluation of CBIR approaches in the context of medical imaging.
Keywords Content-based image retrieval  $\cdot$ Medical imaging  $\cdot$ Pre-trained embeddings  $\cdot$ Re-ranking
1 Introduction
In the realm of computer vision, content-based image retrieval (CBIR) has been the subject of extensive research for several decades (Dubey, 2021). CBIR systems typically utilize low-dimensional image representations stored in a database and subsequently retrieve similar images based on distance metrics or similarity measures of the image representations. Early approaches to CBIR involved manually crafting distinctive features, which led to a semantic gap, resulting in the loss of crucial image details due to the limitations of low-dimensional feature design (Dubey, 2021; Wang et al., 2022). However, recent studies in deep learning have redirected attention towards the creation of machine-generated discriminative feature spaces, effectively addressing and bridging this semantic gap (Qayyum et al., 2017). This shift has significantly enhanced the potential for more accurate and efficient CBIR methods (Dubey, 2021).
While natural image retrieval has been extensively researched, the application of retrieval frameworks to medical images, particularly radiology images, presents ongoing challenges. CBIR offers numerous advantages for medical images. Radiologists can utilize CBIR to search for similar cases, enabling them to review the history, reports, patient diagnoses, and prognoses, thereby enhancing their decision-making process. In real-world use-cases, we often encounter huge anonymized and unannotated datasets available from different studies or institutions where the available meta-information, such as DICOM header data, has been removed or is inconsistent. Manually, searching for relevant images in such databases is extremely time-consuming. Moreover, the development of new tools and research in the medical field requires trustable dataset sources and therefore a reliable method for retrieving images, making CBIR an essential component in advancing computer-aided medical image analysis and diagnosis. One of the key challenges with applying standard CBIR techniques to medical images lies in the fact that algorithms developed for natural images are typically designed for 2D images, while medical images are often 3D volumes which adds a layer of complexity to the retrieval process.
Recent studies have proposed and demonstrated the potential use of pre-trained vision embeddings for CBIR in the context of radiology image retrieval (Khun Jush et al., 2023; Abacha et al., 2023; Denner et al., 2024; Truong et al., 2023). However, these studies have primarily focused on 2D images (Denner et al., 2024) or specific pathologies or tasks (Abacha et al., 2023; Khun Jush et al., 2023; Truong et al., 2023), overlooking the presence of multiple organs in the volumetric images, which is a critical aspect of real-world scenarios. Large multi-organ medical image datasets can be leveraged to thoroughly evaluate the efficacy of the proposed methods, enabling a more comprehensive assessment of CBIR approaches for radiology images. Despite previous efforts, there is still no established benchmark available for comparing methods for the retrieval of 3D volumetric medical images. This absence of a benchmark impedes the ability to objectively evaluate and compare the efficiency of the proposed CBIR approaches in the context of medical imaging.
Our previous work (Khun Jush et al., 2023) demonstrated the potential of utilizing pre-trained embeddings, originally trained on natural images, for various medical image retrieval tasks using the Medical Segmentation Decathlon Challenge (MSD) dataset (Antonelli et al., 2022). The approach is outlined in Figure 1. Building upon this, the current study extends the methodology proposed in Khun Jush et al. (2023) to establish a benchmark for anatomical region-based and localized multi-organ retrieval. While the focus of Khun Jush et al. (2023) was on evaluating the feasibility of using 2D embeddings and benchmarking different aggregation strategies of 2D information for 3D medical image retrieval within the context of the single-organ MSD dataset (Antonelli et al., 2022), it was observed that the single-organ labeling, hinders the evaluations for images containing multiple organs. The main objective of this study is to set a benchmark for organ retrieval at the localized level, which is particularly valuable in practical scenarios, such as when users zoom in on specific regions of interest to retrieve similar images of the precise organ under examination. To achieve this, we evaluate a count-based method in regions using the TotalSegmentator dataset (TS) (Wasserthal et al., 2023). TS dataset along with its detailed multi-organ annotations is a valuable resource for medical image analysis and research. This dataset provides comprehensive annotations for 104 organs or anatomical structures, which allow us to derive fine-grained retrieval tasks and comprehensively evaluate the proposed methods.
The contribution of this work is as follows:
  • We benchmarked pre-trained 2D embeddings trained supervised on medical images against self-supervised pre-trained embeddings trained on non-medical images for 3D radiology image retrieval. We utilize a count-based method to aggregate search results based on slice similarity to volume-level data retrieval.
  • We propose evaluation schemes based on the TotalSegmentator dataset Wang et al. (2022) for 29 aggregated coarse anatomical regions and all 104 original anatomical regions. Our proposed evaluation assesses the capabilities of a 3D image search system at different levels, including a fine-grained measure related to the localization of anatomical regions.
  • We adopted a late interaction re-ranking method originally used for text retrieval called ColBERT (Khattab and Zaharia, 2020) for volumetric image retrieval. For a 3D image query, this two-stage method generates a candidate 3D image result list utilizing a fast slice-wise similarity search and count-based aggregation. In the second stage, the full similarity information between all query and candidate slices is aggregated to determine re-ranking scores.
  • We benchmarked the proposed re-ranking method against the original method proposed in Khun Jush et al. (2023) for volume, region, and localized retrieval on 29 modified coarse anatomical regions and 104 original anatomical regions from TS dataset Wang et al. (2022).
Figure 1: Overview of a retrieval system based on Khun Jush et al. (2023): Step 1: 2D slices are extracted from the 3D volumes. Step 2: Feature extractors are used to extract the embeddings from the database slices and query volumes. Step 3: Database embeddings are indexed using HNSW or LSH indexing. Step 4: Search and slice retrieval is performed, and a hit-table is saved (the hit-table shows the occurrence of volume-ids per each query volume or region saved along with the sum of its total score). Step 5: The results from slice retrieval are aggregated to retrieve the final volume.
2 Materials and Methods
2.1 Vector Database and Indexing
In the context of image search a database typically constitutes the central location where all the representations of the images, a.k.a. embeddings, and their metadata including annotations are stored. A query allows the user or the system to request specific images in various ways, e.g., by inputting a reference image or a textual description. The goal is to search the database for similar images that match the query. Similarly, in this study, the search process entails comparing a query image with images in the database to identify the most similar image using the cosine similarity of the embeddings. Throughout this process, we do not depend on any metadata information at any stage. Metadata-independence is an intended design choice and in contrast to widely used metadata-based image retrieval solutions that frequently lack the necessary specificity in real-world retrieval applications. In small sets, the similarity search is easy but with the growing size of the database, the complexity increases. Accuracy and speed are the key factors in search, thus, naive approaches typically fail in huge datasets.
Indexing in the context of content-based image search involves creating a structured system that allows for efficient storage and retrieval of images based on their visual content. A flat index is the simplest form of indexing, where no modification is made to the vectors before they are used for search. In flat indexing, the query vector is compared to every other full-size vector in the database and their distances are calculated. The nearest k of the searched spaces is then returned as the k-nearest neighbors (kNN). While this method is the most accurate, it comes at the cost of significant search time (Aumüller et al., 2020). To improve search time, two approaches can be employed: reducing the vector size through dimensionality reduction, e.g., by reducing the number of bits representing each vector, or reducing the search scope by clustering or organizing vectors into tree structures based on similarity or distance. This results in the identification of an approximation of the true nearest neighbors, known as approximate nearest neighbor search (ANN) (Aumüller et al., 2020).
There are several ANN methods available. In the context of content-based volumetric medical image retrieval, Khun Jush et al. (2023) compared Locality Sensitive Hashing (LSH) Charikar (2002) and Hierarchical Navigable Small World (HNSW) Malkov and Yashunin (2018) for indexing and search. LSH hashes data points in a way that similar data points are mapped to the same buckets with higher probabilities. This allows for a more efficient search for nearest neighbors by reducing the number of candidates to be examined. HNSW (Malkov and Yashunin, 2018) indexing organizes data into a hierarchical graph structure where each layer of the hierarchy has a lower resolution. The top layer connects data points directly, but the lower layers have fewer connections. The graph structure is designed to allow for efficient navigation during the search. Compared to LSH, HNSW typically enables faster search and requires less memory Taha et al. (2024). Based on findings in (Khun Jush et al., 2023) HSNW was chosen as the indexing method in the setting of this study due to speed advantages over LSH at a comparable recall. There are various index solutions available to store and search vectors. In this study, we used the Facebook AI Similarity Search (FAISS) package that enables fast similarity search (Johnson et al., 2019). The indexing process involves running the feature extractors on slices of each volumetric image and storing the output embeddings per slice. The produced representations are then added to the search index which is used later on for vector-similarity-based retrieval.
2.2 Feature Extractors
We extend the analysis of Khun Jush et al. (2023) by adding two ResNet50 embeddings and evaluating the performance of six different slice embedding extractors for CBIR tasks. All the feature extractors are based on deep-learning-based models.
Self-supervised Models: We employed three self-supervised models pre-trained on ImageNet (Deng et al., 2009). DINOv1 (Caron et al., 2021), that demonstrated learning efficient image representations from unlabeled data using self-distillation. DINOv2 (Oquab et al., 2023), is built upon DINOv1 (Caron et al., 2021), and this model scales the pre-training process by combining an improved training dataset, patchwise objectives during training and introducing a new regularization technique, which gives rise to superior performance on segmentation tasks. DreamSim (Fu et al., 2023), built upon the foundation of DINOv1 (Caron et al., 2021), fine-tunes the model using synthetic data triplets specifically designed to be cognitively impenetrable with human judgments. For the self-supervised models, we used the best-performing backbone reported by the developers of the models.
Supervised Models: We included a SwinTransformer model (Liu et al., 2021) and a ResNet50 model (He et al., 2016) trained in a supervised manner using the RadImageNet dataset (Mei et al., 2022) that includes 5 million annotated 2D CT, MRI, and ultrasound images of musculoskeletal, neurologic, oncologic, gastrointestinal, endocrine, and pulmonary pathology. Furthermore, a ResNet50 model pre-trained on rendered images of fractal geometries was included based on (Kataoka et al., 2022). These training images are formula-derived, non-natural, and do not require any human annotation.
2.3 Dataset and Pre-processing
Table 1: Mapping of the original TS classes to 29 coarse anatomical regions.
Anatomical region Mapped class Anatomical region Mapped class Anatomical region Mapped class
adrenal gland left adrenal gland iliopsoas right iliopsoas rib right 11 rib
adrenal gland right adrenal gland inferior vena cava cardiovascular system rib right 12 rib
aorta cardiovascular system kidney left kidney sacrum sacrum
autochthon left autochthon kidney right kidney scapula left scapula
autochthon right autochthon liver liver scapula right scapula
brain brain lung lower lobe left lung small bowel small bowel
clavicula left clavicula lung lower lobe right lung spleen spleen
clavicula right clavicula lung middle lobe right lung stomach stomach
colon colon lung upper lobe left lung trachea trachea
duodenum duodenum lung upper lobe right lung urinary bladder urinary bladder
esophagus esophagus pancreas pancreas vertebrae C1 vertebrae
face face portal and splenic vein portal & splenic vein vertebrae C2 vertebrae
femur left femur pulmonary artery cardiovascular system vertebrae C3 vertebrae
femur right femur rib left 1 rib vertebrae C4 vertebrae
gallbladder gallbladder rib left 2 rib vertebrae C5 vertebrae
gluteus maximus left gluteus muscles rib left 3 rib vertebrae C6 vertebrae
gluteus maximus right gluteus muscles rib left 4 rib vertebrae C7 vertebrae
gluteus medius left gluteus muscles rib left 5 rib vertebrae L1 vertebrae
gluteus medius right gluteus muscles rib left 6 rib vertebrae L2 vertebrae
gluteus minimus left gluteus muscles rib left 7 rib vertebrae L3 vertebrae
gluteus minimus right gluteus muscles rib left 8 rib vertebrae L4 vertebrae
heart atrium left cardiovascular system rib left 9 rib vertebrae L5 vertebrae
heart atrium right cardiovascular system rib left 10 rib vertebrae T1 vertebrae
heart myocardium cardiovascular system rib left 11 rib vertebrae T2 vertebrae
heart ventricle left cardiovascular system rib left 12 rib vertebrae T3 vertebrae
heart ventricle right cardiovascular system rib right 1 rib vertebrae T4 vertebrae
hip left hip rib right 2 rib vertebrae T5 vertebrae
hip right hip rib right 3 rib vertebrae T6 vertebrae
humerus left humerus rib right 4 rib vertebrae T7 vertebrae
humerus right humerus rib right 5 rib vertebrae T8 vertebrae
iliac artery left cardiovascular system rib right 6 rib vertebrae T9 vertebrae
iliac artery right cardiovascular system rib right 7 rib vertebrae T10 vertebrae
iliac vena left cardiovascular system rib right 8 rib vertebrae T11 vertebrae
iliac vena right cardiovascular system rib right 9 rib vertebrae T12 vertebrae
iliopsoas left iliopsoas rib right 10 rib
We designed a CBIR benchmark relying on the publicly available TotalSegmentator (TS) dataset Wasserthal et al. (2023), version 1. This dataset comprises in total of 1204 computed tomography (CT) volumes covering 104 anatomical structure annotations (TS, V1). The anatomical regions presented in the original dataset include several fine-grained sub-classes for which we considered an aggregation to a coarser common class as a reasonable measure, e.g., all the rib classes are mapped to a single class ‘rib’. The coarse organ labels can help identify similarities and potential mismatches between neighboring anatomical regions, providing valuable insights into the proximity information of the target organ. LABEL:tab:_mapping_table shows the mapping of the original TS classes to the coarse aggregated classes. For the sake of reproducibility, the query cases are sourced from the original TS test split, while the cases contained in the original TS train and validation set serve as the database for searching. The search is assessed on the retrieval rate of 29 coarse anatomical structures and 104 original TS anatomical structures.
The models presented in Section 2.2 are 2D models used without fine-tuning to extract the embeddings. Thus, per each 3D volume, individual 2D slices of the corresponding 3D volumes are utilized for embedding extraction. The input size for all the used models is equal to $224\times 224$ pixels with image replication along the RGB channel axis. For all the ViT-based models and the ResNet50 trained on fractal images, images are normalized to the ImageNet mean and standard deviation of (.485,.456,.406).485.456.406(.485,.456,.406)( .485 , .456 , .406 ) and (.229,.224,.225).229.224.225(.229,.224,.225)( .229 , .224 , .225 ), respectively. For the SwinTransformer and the ResNet50 model pre-trained on the RadImageNet dataset, the images are normalized to .5.5.5.5 mean and .5.5.5.5 standard deviation based on Mei et al. (2022). The total size of the database is $290757$ embeddings, while the final query set of the test set comprises $20442$ embeddings.
2.4 Search and Retrieval
After creating the vector database, the search is performed using the embeddings extracted from slices of query volumes. The simplest way of retrieval is to match a 2D query slice $q$ with the most similar 2D slice in the database $s^{*}$ by finding the slice-embedding that maximizes the cosine similarity with respect to the embedding associated with $q$, i.e.
$s^{*}=\underset{s\in\text{Data}}{\mathrm{argmax}}\frac{\langle\phi(s),\phi(q)\rangle}{\|\phi(s)\|_{2}\|\phi(q)\|_{2}}=\underset{s\in\text{Data}}{\mathrm{argmax}}\left\langle v_{s},\frac{\phi(q)}{\|\phi(q)\|_{2}}\right\rangle$ (1)
[…]
2.4.1 Volume-based retrieval and evaluation
Figure 2: Volume-based retrieval: For a query volume $V_{q}$ covering a range of anatomical regions, a volume is retrieved that should cover the same anatomical regions. The similarity search is based on all slices from the query volume.
In this setting, we want to assess the capability of the search system to retrieve an image volume that covers the same anatomical regions overall as a given query volume $V_{Q}=[q_{1},...,q_{n}]$. For every slice $q_{i}$ from the query volume $V_{Q}$, the system retrieves the most similar slice $s_{i}^{*}$ from the database using (1). Subsequently, the corresponding volume-id and its similarity score are stored in a hit-table similar to the hit-table shown in Figure 1. We then apply the count-based aggregation scheme from Khun Jush et al. (2023) that uses the hit-table to identify the volume $V_{R}$ with the most hits given the query volume at hand, see Figure 1. The system then returns $V_{R}$ as the final result of the similarity search given the query volume $V_{Q}$. Abacha et al. (2023) refers to this method as the retrieval based on frequency. The evaluation of the search system on the proposed benchmark dataset is done using the measure of Recall (True Positive Rate). To this end, the aggregated labels of the query volumes are compared against the aggregated labels of the associated retrieved volumes (see 3.1.2). The recall in this setting is high if the aggregated query and retrieved volume labels coincide. An overview is shown in Figure 2.
2.4.2 Region-based retrieval and evaluation
In this setting the search system is queried with an image (sub-)volume which is constrained to a specific anatomical sub-region (e.g. liver, pancreas, heart,…). For each anatomical region, we want to individually assess the capability of the system to retrieve an image volume containing the anatomical region.
The query (sub-)volumes for different anatomical regions are generated as follows. Given a selected anatomical region $r$ and a query image volume $V_{Q}=[q_{1},...,q_{n}]$, the smallest subset slices $V_{Q,r}=[q_{m},...,q_{k}]\subset V_{Q}$ is chosen that entirely contains the anatomical region $r$ visible in $V_{Q}$. Based on the sub-volume $V_{Q,r}$ a similarity search is conducted to build up a hit-table, and the count-based aggregation is conducted to finally retrieve for this query the volume with most hits, as described in Section 2.4.1.
Figure 3: Region-based retrieval. Anatomical regions are considered individually. A sub-volume constrained to an anatomical region of interest $r$ is generated and fed to the search system to retrieve a volume containing the anatomical region. A case is considered a True Positive (TP) if the retrieved case contains the region $r$ at some location.
[…]
2.4.3 Localized retrieval and evaluation
Figure 4: Localized retrieval. Anatomical regions are considered individually. A sub-volume constrained to the anatomical region of interest $r$ is generated and fed to the search system to retrieve a volume containing the same anatomical region. A case is only considered as True Positive (TP) if at least one of the slices in the retrieved volume contains the region $r$.
In this setting, the system is queried with an image sub-volume which is constrained to a specific anatomical sub-region (e.g. liver, pancreas, heart,…). For each anatomical region, we want to individually assess the capability of the system to retrieve an image volume containing the anatomical region and to localize the region of interest within the retrieved volume.
[…]
Another measure to assess the capability of the system to localize a region can be defined as the ratio of the slices that actually contain the subregion $r$ in the retrieved volume to the total number of slices hit in the retrieved volume. In detail, the localization-ratio (LR) is defined as:
LR=|slices hit in retrieved vol. that contain r||slices hit in retrieved $\mathrm{LR}=\frac{|\text{slices hit in retrieved vol. that contain r}|}{|\text{slices hit in retrieved vol.}|}$ (2)
For example, the query consists of $60$ slices related to region $r$. The table representing the top 3 volumes hit count is $[48,21,4]$. In the volume with the top hit count, 12 out of the 48 hit slices actually contain region $r$, indicating successful localization. The localization-ratio is then given by $12/48=.25$.
2.5 Re-ranking retrieval and evaluation
Re-ranking in information retrieval involves the process of re-ordering the initially retrieved results to better align with the user’s information needs. This can be achieved through different methods such as relevance feedback, learning to rank algorithms, or incorporating contextual information (Ai et al., 2018; Guo et al., 2020; MacAvaney et al., 2019). Relevance feedback allows users to provide input on the initial results, which is then used to adjust the ranking (Ai et al., 2018). Learning to rank algorithms utilizes machine learning techniques to re-rank results based on relevant features (Guo et al., 2020). Additionally, re-ranking methods may also consider contextual information such as user behavior, temporal relevance, or other relevant factors to better reflect the user’s current information needs, ultimately enhancing the overall quality of retrieved results (MacAvaney et al., 2019). A method based on contextualized information proposed in Khattab and Zaharia (2020) called ColBERT (Contextualized Late Interaction over BERT). ColBERT operates by generating contextualized representations of the query and the documents using BERT (Devlin et al., 2018). In this method, queries and documents are encoded into more detailed multi-vector representations, and relevance is gauged through comprehensive yet scalable interactions between these sets of vectors. ColBERT creates an embedding for each token in the query and document, and it measures relevance as the total of maximum similarities between each query vector and all vectors within the document (Santhanam et al., 2021). This late interaction approach allows for a more refined and contextually aware retrieval process, thereby enhancing the quality of information retrieval.
Inspired by ColBERT we introduce a two-stage method in which filtering of the search space is performed and the total similarity of the entire target volume is considered to re-rank and score the retrieved volumes. To create an analogy to the ColBERT method each word can be considered as one slice and each passage of the database or each question of the query can be considered as one volume. Instead of the BERT encoder for the image retrieval task, the pre-trained vision models can be used to create the embeddings as discussed in Section 2.2.
Figure 5: Overview of re-ranking. Step 1: Filtering based on at least one similar slice leads to the selection of candidate volumes Step 2: followed by similarity score computation using dot product on the normalized embedding matrices. Step 3: The final step involves max-pooling and summation to determine the top-scoring volumes for retrieval.
An overview of the proposed method is shown in Figure 5. The method consists of the following steps:
2.5.1 Step 1: Filtering
We assume that the embeddings have been computed and stored as vectors for all image data. In the first step, for each slice $q_{i}$ of the query volume $V_{Q}=[q_{1},...,q_{N}]$, the most similar slice $s_{i}^{*}$ from the database is retrieved via similarity search (1). We denote volumes associated to $s_{1}^{*},...,s_{N}^{*}$ by $V_{1},...,V_{N}$. Clearly, for some indices $i,j$ the retrieved sliced can belong to the same volume, i.e. $V_{i}=V_{j}$. We filter out duplicates for our list and let $V_{1},..,V_{M}$ with $M\leq N$ denote the unique volumes associated with the retrieved slices. These unique volumes resulting from this initial search stage are considered as candidates for a more compute-intensive re-ranking procedure which takes all similarity interactions between all the volume slices into account.
2.5.2 Step 2: Similarity Score on Embedding Matrix
We assume that our slice embedding $\phi$ maps each slice to a vector of fixed length $L$. Hence, for any image volume $V=[v_{1},...,v_{n}]$ with $n$ slices $v_{1},...,v_{n}$ we can compute the $L_{2}$-normalized embedding matrix
$M_{V}=\left[\frac{\phi(v_{1})}{\|\phi(v_{1})\|_{2}},...,\frac{\phi(v_{n})}{\|\phi(v_{n})\|_{2}}\right]$ (3)
of dimension $n\times L$. Given another image volume $W=[w_{1},...,w_{m}]$ of size $m\times L$ we can compute the similarity matrix of size $n\times m$ via the matrix dot product:
$\mathrm{SIM}(M_{V},M_{W})=\left[\frac{\langle\phi(v_{i}),\phi(w_{j})\rangle}{\|\phi(v_{i})\|_{2}~{}\|\phi(w_{j})\|_{2}}\right]_{\begin{subarray}{c}i=1,...,n\\ j=1,...,m\end{subarray}}$ (4)
The entry $(i,j)$ of this matrix contains the cosine-similarity score of the embeddings related to slice $i$ of volume $V$ and slice $j$ of volume $W$. For our re-ranking, we first compute in this step, the embedding matrix $M_{V_{Q}}$ for the query volume $V_{Q}$ and the embedding matrices $M_{V_{1}},...,M_{V_{M}}$ for all unique retrieved volumes $V_{1},...,V_{M}$ from Step 1. Then we compute the similarity matrices $\mathrm{SIM}(M_{V_{Q}},M_{V_{k}})$ for $k=1,...,M$.
2.5.3 Step 3: Final Score and Re-ranking
To compute the rank score of each volume $V_{i}$, the associated similarity matrix $\mathrm{SIM}(M_{V_{Q}},M_{V_{i}})$ is max-pooled row-wise which corresponds to determine the slice in $V_{i}$ with highest cosine similarity to a given slice in $V_{Q}$. The resulting vector of length $n$ is then summed to obtain the total maximum slice similarity as final rank score (RS), i.e. for $k=1,...,M$ we compute:
$\mathrm{RS}(V_{k})=\sum_{i=1}^{n}\max_{j=1,...,m_{k}}\mathrm{SIM}(M_{V_{Q}},M_{V_{k}})_{i,j}$ (5)
where ${V_{k}}_{j}$ denotes the $j$-th slice of volume $V_{k}$ and $m_{k}$ is the total number of slices of ${V_{k}}$. The retrieved volumes are re-ranked according to their RS, i.e. the volume with the highest score is considered as the most relevant search result.
2.5.4 Localization with Re-ranking
In section 2.4.3 we proposed a measure for assessing the capabilities of the search system to localize anatomical regions of interest. This localization measure relied on the hit slices that were finally retrieved by count-based aggregation. As described above, re-ranking may select a volume with a lower total hit count as the final search result based on a more fine-grained similarity measure. We propose to utilize this similarity information to identify slices of maximal similarity with the query data in order to localize the anatomical region of interest.
In detail, after having obtained the most relevant search result via re-ranking, we need to identify the most relevant slices that could contain the anatomical region of interest $r$. To this end, we consider the slices that highly contribute to the final RS score (5). Since the full similarity matrices (4) for all candidates from the first search stage are available, we can utilize these to identify the locations of maximum similarity between the query and candidate slices.
Assume that the volume $V_{k^{\star}}$ obtained the highest RS score in the re-ranking process and therefore was returned as the final result after re-ranking, i.e. $k^{*}=argmax_{k}(RS(V_{k}))$. Further assume that $V_{k^{\star}}$ consists of $m_{k^{\star}}$ slices. First, we compute the vector $m_{\mathrm{SIM}}$ of size $m_{k^{\star}}$ containing for each slice of $V_{k^{\star}}$ the maximum similarity with the query slices:
$m_{\mathrm{SIM}}(V_{k^{\star}})=\left[\max_{i=1,...,n}\mathrm{SIM}(M_{V_{Q}},M_{V_{k^{\star}}})_{i,j}\right]_{j=1,...,m_{k^{\star}}}$ (6)
Hence, to determine the $L\leq m_{k^{\star}}$ slices of $V_{k^{\star}}$ with the maximum similarity, we determine the indices of the $L$ largest values of the vector $m_{\mathrm{SIM}}$. These slices are considered candidate locations for the anatomical region $r$. In our experiments, we chose $L=15$. The localization-ratio (2) can be adopted in the case of re-ranking by determining how many of the $L$ most similar slices actually contained the anatomical region of interest.
LR=|slices among the L most similar slices of Vk that contain $\mathrm{LR}=\frac{|\text{slices among the $L$ most similar slices of $V_{k^{\star}}$ that contain r}|}{L}$ (7)
3 Evaluation
In this section, we evaluate the retrieval recall of the methods explained in Section 2.4 and Section 2.5. The results related to the 29 coarse anatomical structures from LABEL:tab:_mapping_table and the results related to the original 104 fine-grained anatomical structures fromWasserthal et al. (2023) are presented separately in the following. In the tables presented in this section, the average and standard deviation (STD) columns allow identifying difficult classes across models (low average) and the ones that have higher variations among models (higher STD). The average and STD rows show the average and STD over all the classes for each model.
3.1 Search and Retrieval
3.1.1 Slice-wise
Detailed computation of the recall measure for different retrieval methods is explained in Section 2.4. LABEL:tab:slicewise-recall-28-regions and LABEL:tab:slicewise-recall-104-regions show the retrieval recall of 29 coarse anatomical regions and 104 original TS anatomical regions, respectively, using the slice-wise method. The slice-wise recall is considered the lower bound recall because for a perfect recall all the anatomical regions present in the query slice should appear in the retrieved slice.
In slice-wise retrieval, DreamSim is the best-performing model with retrieval recall of .$863\pm.107$ and .$9.7$129 for coarse and original TS classes, respectively. ResNet50 pre-trained on fractal images has the lowest retrieval recall almost on every anatomical region for 29 and 104 classes. This is however expected due to the nature of synthetic generated images.
In LABEL:tab:slicewise-recall-104-regions the gallbladder has the lowest retrieval rate followed by vertebrae C4 and C5 (see average column). However, in LABEL:tab:slicewise-recall-28-regions the vertebrae class shows a higher recall which indicated that the vertebrae classes were detected but the exact location, i.e. C4 or C5 were mismatched. The same pattern can be observed in rib classes.
Table 2: Slice-wise recall of coarse anatomical regions (29 classes) using HNSW Indexing. In each row, bold numbers represent the best-performing values, while italicized numbers indicate the worst-performing. The separate average and standard deviation (STD) columns are color-coded, with blue indicating the best-performing values and yellow indicating the worst-performing values across different models. Additionally, bold numbers in colored columns represent the best classes in terms of average and standard deviation, while italicized values represent the worst-performing class across the models.
Model DINOv1 DINOv2 DreamSim SwinTrans. ResNet50
Dataset (pre-trained) (ImgNet) (ImgNet) (ImgNet) (RadImg) (Fractaldb) (RadImg) Average STD
adrenal gland .749 .639 .671 .614 .490 .557 .620 .090
autochthon .980 .974 .979 .976 .941 .965 .969 .015
brain .852 .843 .901 .894 .850 .863 .867 .025
cardiovascular system .978 .974 .979 .970 .941 .953 .966 .015
clavicula .886 .884 .898 .857 .632 .873 .838 .102
colon .932 .931 .945 .912 .830 .905 .909 .042
duodenum .678 .682 .719 .697 .605 .733 .686 .045
esophagus .934 .934 .936 .933 .870 .894 .917 .028
face .854 .840 .872 .788 .692 .733 .797 .072
femur .927 .907 .953 .914 .778 .860 .890 .063
gallbladder .246 .345 .312 .341 .347 .400 .332 .051
gluteus muscles .964 .940 .978 .950 .879 .915 .938 .036
hip .959 .928 .974 .941 .880 .907 .931 .034
humerus .575 .600 .633 .598 .351 .523 .547 .102
iliopsoas .950 .933 .957 .934 .863 .923 .927 .034
kidney .759 .771 .791 .776 .641 .776 .752 .055
liver .840 .817 .844 .841 .814 .839 .833 .013
lung .953 .930 .958 .940 .890 .898 .928 .028
pancreas .720 .685 .779 .734 .552 .722 .699 .078
portal and splenic vein .731 .627 .679 .658 .522 .584 .634 .074
rib .950 .942 .951 .948 .900 .933 .937 .020
sacrum .894 .865 .907 .878 .805 .856 .867 .036
scapula .935 .913 .924 .891 .793 .869 .887 .052
small bowel .896 .872 .900 .894 .783 .892 .873 .045
spleen .774 .719 .735 .699 .731 .693 .725 .029
stomach .811 .781 .844 .778 .741 .752 .784 .038
trachea .893 .862 .903 .863 .762 .816 .850 .053
urinary bladder .720 .643 .722 .720 .633 .666 .684 .041
vertebrae .981 .967 .977 .969 .950 .964 .968 .011
Average .855 .832 .863 .837 .751 .813
STD .108 .118 .107 .114 .152 .124
Table 3: Slice-wise recall of all TS anatomical regions (104 classes) using HNSW Indexing. In each row, bold numbers represent the best-performing values, while italicized numbers indicate the worst-performing. The separate average and standard deviation (STD) columns are color-coded, with blue indicating the best-performing values and yellow indicating the worst-performing values across different models. Additionally, bold numbers in colored columns represent the best classes in terms of average and standard deviation, while italicized values represent the worst-performing class across the models.
Model DINOv1 DINOv2 DreamSim SwinTrans. ResNet50
Dataset (pre-trained) (ImgNet) (ImgNet) (ImgNet) (RadImg) (Fractaldb) (RadImg) Average STD
adrenal gland left .636 .524 .573 .539 .407 .453 .522 .082
adrenal gland right .644 .515 .593 .551 .408 .521 .539 .080
aorta .954 .941 .946 .952 .915 .926 .939 .015
autochthon left .981 .972 .980 .974 .942 .966 .969 .014
autochthon right .980 .974 .979 .976 .942 .965 .969 .014
brain .852 .843 .901 .894 .850 .863 .867 .025
clavicula left .866 .875 .886 .864 .636 .874 .833 .097
clavicula right .862 .871 .867 .840 .614 .855 .818 .101
colon .932 .931 .945 .912 .830 .905 .909 .042
duodenum .678 .682 .719 .697 .605 .733 .686 .045
esophagus .934 .934 .936 .933 .870 .894 .917 .028
face .854 .840 .872 .788 .692 .733 .797 .072
femur left .920 .902 .940 .909 .773 .855 .883 .061
femur right .931 .910 .952 .938 .808 .915 .909 .052
gallbladder .246 .345 .312 .341 .347 .400 .332 .051
gluteus maximus left .937 .914 .951 .927 .845 .903 .913 .037
gluteus maximus right .942 .914 .945 .925 .858 .900 .914 .032
gluteus medius left .930 .878 .948 .920 .824 .883 .897 .045
gluteus medius right .922 .892 .951 .923 .852 .893 .905 .034
gluteus minimus left .872 .824 .894 .855 .795 .876 .853 .037
gluteus minimus right .876 .811 .878 .874 .819 .898 .860 .035
heart atrium left .709 .656 .800 .680 .588 .542 .663 .091
heart atrium right .793 .762 .870 .773 .684 .668 .758 .074
heart myocardium .798 .757 .844 .808 .715 .733 .776 .049
heart ventricle left .778 .724 .824 .788 .699 .720 .756 .048
heart ventricle right .802 .801 .851 .822 .723 .738 .789 .049
hip left .959 .928 .971 .937 .880 .905 .930 .034
hip right .963 .932 .977 .948 .889 .916 .938 .032
humerus left .525 .571 .591 .577 .313 .471 .508 .105
humerus right .593 .625 .627 .567 .314 .529 .543 .118
iliac artery left .882 .863 .902 .893 .813 .841 .866 .034
iliac artery right .905 .869 .918 .895 .822 .851 .876 .036
iliac vena left .903 .868 .908 .893 .825 .857 .876 .032
iliac vena right .910 .870 .923 .891 .831 .873 .883 .033
iliopsoas left .950 .929 .958 .932 .861 .924 .926 .034
iliopsoas right .947 .929 .951 .932 .854 .922 .923 .035
inferior vena cava .928 .896 .922 .923 .841 .893 .901 .033
kidney left .719 .708 .762 .747 .600 .762 .716 .061
kidney right .708 .724 .755 .737 .602 .756 .714 .058
liver .840 .817 .844 .841 .814 .839 .833 .013
lung lower lobe left .903 .885 .908 .887 .826 .811 .870 .041
lung lower lobe right .903 .880 .914 .897 .806 .809 .868 .048
lung middle lobe right .800 .785 .818 .794 .726 .699 .770 .047
lung upper lobe left .917 .909 .921 .906 .850 .875 .896 .028
lung upper lobe right .928 .883 .919 .885 .818 .848 .880 .042
pancreas .720 .685 .779 .734 .552 .722 .699 .078
portal and splenic vein .731 .627 .679 .658 .522 .584 .634 .074
pulmonary artery .819 .711 .773 .679 .526 .563 .679 .115
rib left 1 .855 .824 .867 .851 .669 .821 .815 .073
rib left 10 .827 .775 .803 .823 .742 .747 .786 .037
rib left 11 .773 .767 .785 .788 .694 .756 .761 .034
rib left 12 .594 .568 .682 .620 .481 .576 .587 .066
rib left 2 .841 .804 .858 .807 .681 .803 .799 .062
rib left 3 .832 .803 .808 .805 .728 .789 .794 .035
rib left 4 .820 .783 .809 .776 .738 .759 .781 .031
rib left 5 .789 .786 .805 .784 .699 .723 .764 .043
rib left 6 .815 .787 .797 .787 .706 .751 .774 .039
rib left 7 .830 .825 .834 .829 .734 .778 .805 .040
rib left 8 .810 .799 .850 .831 .745 .777 .802 .038
rib left 9 .826 .803 .833 .853 .737 .780 .805 .042
rib right 1 .852 .820 .828 .831 .672 .827 .805 .066
rib right 10 .827 .768 .804 .814 .728 .747 .781 .040
rib right 11 .770 .763 .798 .771 .681 .742 .754 .040
rib right 12 .577 .570 .619 .634 .456 .556 .569 .063
rib right 2 .839 .820 .840 .815 .680 .802 .799 .060
rib right 3 .850 .794 .826 .795 .725 .780 .795 .043
rib right 4 .834 .790 .809 .770 .738 .753 .782 .036
rib right 5 .802 .791 .810 .776 .709 .718 .768 .044
rib right 6 .810 .788 .772 .779 .709 .741 .766 .036
rib right 7 .803 .813 .805 .817 .731 .765 .789 .034
rib right 8 .814 .792 .847 .833 .754 .778 .803 .035
rib right 9 .823 .793 .813 .844 .738 .776 .798 .038
sacrum .894 .865 .907 .878 .805 .856 .867 .036
scapula left .922 .891 .908 .891 .798 .884 .882 .044
scapula right .930 .905 .919 .884 .799 .872 .885 .047
small bowel .896 .872 .900 .894 .783 .892 .873 .045
spleen .774 .719 .735 .699 .731 .693 .725 .029
stomach .811 .781 .844 .778 .741 .752 .784 .038
trachea .893 .862 .903 .863 .762 .816 .850 .053
urinary bladder .720 .643 .722 .720 .633 .666 .684 .041
vertebrae C1 .555 .571 .655 .592 .399 .592 .561 .086
vertebrae C2 .744 .613 .812 .594 .529 .643 .656 .104
vertebrae C3 .677 .566 .586 .414 .271 .359 .479 .155
vertebrae C4 .427 .377 .519 .488 .323 .308 .407 .087
vertebrae C5 .513 .444 .572 .565 .330 .366 .465 .102
vertebrae C6 .562 .562 .536 .423 .220 .445 .458 .131
vertebrae C7 .712 .645 .685 .576 .375 .580 .595 .121
vertebrae L1 .620 .561 .662 .653 .452 .540 .582 .080
vertebrae L2 .555 .514 .658 .587 .411 .591 .553 .084
vertebrae L3 .747 .533 .608 .593 .503 .629 .602 .086
vertebrae L4 .612 .449 .639 .693 .523 .572 .581 .087
vertebrae L5 .732 .592 .748 .714 .606 .631 .670 .069
vertebrae T1 .723 .698 .704 .689 .464 .694 .662 .098
vertebrae T10 .568 .517 .583 .565 .449 .518 .533 .050
vertebrae T11 .555 .539 .574 .546 .433 .506 .526 .051
vertebrae T12 .556 .554 .604 .581 .468 .562 .554 .047
vertebrae T2 .750 .670 .752 .726 .563 .679 .690 .071
vertebrae T3 .794 .744 .814 .736 .621 .668 .729 .073
vertebrae T4 .742 .715 .713 .666 .540 .666 .674 .072
vertebrae T5 .647 .618 .701 .627 .513 .550 .609 .068
vertebrae T6 .696 .627 .637 .597 .488 .514 .593 .079
vertebrae T7 .703 .680 .705 .613 .460 .516 .613 .104
vertebrae T8 .595 .590 .728 .564 .450 .469 .566 .101
vertebrae T9 .603 .540 .660 .609 .515 .524 .575 .057
Average .784 .750 .797 .765 .659 .726
STD .137 .144 .129 .140 .172 .154
3.1.2 Volume-based
This section presents the recall of volume-based retrieval explained in Section 2.4.1 An overview of the evaluation is shown in Figure 2. In volume-based retrieval, per each query volume, one volume is retrieved. In the recall computation, the classes present in both the query and the retrieved volume are considered TP classes. The classes that are present in the query volume and are missing from the retrieved volume are considered FN.
LABEL:tab:volumewise-recall-29-regions and LABEL:tab:volumewise-recall-104-regions present the retrieval recall of the volume-based method on 29 and 104 classes, respectively. The overall recall rates are increased compared to slice-wise retrieval which is expected due to the aggregation and contextual effects of neighboring slices.
LABEL:tab:volumewise-recall-29-regions shows that ResNet50 trained on RadImageNet outperforms other methods with an average recall of .$952\pm.043$. However, in LABEL:tab:volumewise-recall-104-regions DINOv1 outperforms all models including ResNet50 with an average recall of .$923\pm.077$. This shows that the embeddings of finer classes are retrieved and assigned to a different similar class by ResNet50, thus, the performance from fine to coarse classes is improved. Whereas, all the self-supervised methods in LABEL:tab:volumewise-recall-104-regions outperform the supervised methods. Although some models perform slightly better than others based on looking at isolated classes, overall models perform on par.
Table 4: Volume-based retrieval recall of coarse anatomical regions (29 classes) using HNSW Indexing. In each row, bold numbers represent the best-performing values, while italicized numbers indicate the worst-performing. The separate average and standard deviation (STD) columns are color-coded, with blue indicating the best-performing values and yellow indicating the worst-performing values across different models. Additionally, bold numbers in colored columns represent the best classes in terms of average and standard deviation, while italicized values represent the worst-performing class across the models.
Model DINOv1 DINOv2 DreamSim SwinTrans. ResNet50
Dataset (pre-trained) (ImgNet) (ImgNet) (ImgNet) (RadImg) (Fractaldb) (RadImg) Average STD
adrenal gland 1.000 .960 .960 .960 .980 .980 .973 .016
autochthon .985 .969 .969 .985 .985 1.000 .982 .012
brain .692 .769 .769 .846 .769 .846 .782 .058
cardiovascular system 1.000 1.000 .968 1.000 1.000 .968 .989 .016
clavicula .949 .949 .949 .897 .821 .949 .919 .052
colon 1.000 .943 .981 .943 1.000 .962 .972 .026
duodenum .940 .860 .900 .920 .980 .920 .920 .040
esophagus .964 .964 .946 .982 .946 1.000 .967 .021
face .765 .765 .706 .706 .706 .882 .755 .069
femur .933 .933 .911 .911 .956 .933 .930 .017
gallbladder .846 .795 .872 .821 .846 .897 .846 .036
gluteus muscles 1.000 .977 .977 .977 1.000 .955 .981 .017
hip 1.000 .977 .977 .977 1.000 .955 .981 .017
humerus .898 .857 .980 .898 .878 .878 .898 .043
iliopsoas .981 .981 .962 .962 .981 .962 .972 .010
kidney .945 .927 .945 .891 .927 .964 .933 .025
liver .964 .945 .982 .945 .982 .982 .967 .018
lung .983 .983 .931 .983 .983 .983 .974 .021
pancreas .940 .920 .920 .940 .960 .980 .943 .023
portal and splenic vein .980 .960 .940 .980 .960 .980 .967 .016
rib .983 .983 .949 1.000 .966 1.000 .980 .020
sacrum .977 .955 .977 .955 1.000 .955 .970 .019
scapula .909 .909 .909 .818 .886 .886 .886 .035
small bowel .958 .896 .917 .958 .979 .938 .941 .031
spleen 1.000 .980 .960 .980 .980 1.000 .983 .015
stomach 1.000 .980 .961 .980 .980 1.000 .984 .015
trachea .951 .951 .951 .878 .805 .902 .907 .059
urinary bladder 1.000 .977 .977 .953 .977 .953 .973 .018
vertebrae .984 .969 1.000 .984 1.000 1.000 .990 .013
Average .949 .932 .936 .932 .939 .952
STD .072 .064 .063 .067 .078 .043
Table 5: Volume-based retrieval recall of all TS anatomical regions (104 classes) using HNSW Indexing. In each row, bold numbers represent the best-performing values, while italicized numbers indicate the worst-performing. The separate average and standard deviation (STD) columns are color-coded, with blue indicating the best-performing values and yellow indicating the worst-performing values across different models. Additionally, bold numbers in colored columns represent the best classes in terms of average and standard deviation, while italicized values represent the worst-performing class across the models.
Model DINOv1 DINOv2 DreamSim SwinTrans. ResNet50
Dataset (pre-trained) (ImgNet) (ImgNet) (ImgNet) (RadImg) (Fractaldb) (RadImg) Average STD
adrenal gland left .960 .900 .920 .900 .900 .960 .923 .029
adrenal gland right .980 .900 .940 .900 .900 .960 .930 .035
aorta .984 .934 .934 .934 .918 .934 .940 .022
autochthon left .969 .923 .938 .923 .908 .954 .936 .023
autochthon right .969 .923 .938 .923 .908 .954 .936 .023
brain .692 .692 .692 .692 .769 .692 .705 .031
clavicula left .949 .923 .897 .821 .821 .897 .885 .053
clavicula right .974 .947 .921 .816 .816 .895 .895 .067
colon .981 .906 .962 .906 .906 .925 .931 .033
duodenum .920 .820 .880 .880 .880 .880 .877 .032
esophagus .946 .911 .911 .911 .875 .964 .920 .031
face .765 .706 .647 .588 .706 .765 .696 .069
femur left .911 .911 .889 .867 .867 .911 .893 .022
femur right .927 .927 .902 .902 .927 .927 .919 .013
gallbladder .846 .744 .846 .795 .795 .872 .816 .047
gluteus maximus left .977 .953 .953 .907 .930 .930 .942 .024
gluteus maximus right .977 .953 .930 .930 .907 .930 .938 .024
gluteus medius left .977 .932 .955 .909 .909 .932 .936 .027
gluteus medius right .977 .930 .953 .930 .907 .953 .942 .024
gluteus minimus left .977 .953 .953 .907 .930 .930 .942 .024
gluteus minimus right .976 .952 .952 .905 .929 .952 .944 .025
heart atrium left .915 .830 .872 .936 .830 .979 .894 .060
heart atrium right .939 .898 .898 .939 .816 .980 .912 .056
heart myocardium .939 .898 .898 .939 .816 .980 .912 .056
heart ventricle left .939 .898 .878 .939 .816 .980 .908 .057
heart ventricle right .939 .898 .898 .939 .816 .980 .912 .056
hip left .977 .932 .955 .932 .909 .932 .939 .023
hip right .977 .932 .955 .932 .886 .932 .936 .030
humerus left .949 .897 .949 .872 .846 .897 .902 .041
humerus right .875 .854 .917 .833 .833 .813 .854 .037
iliac artery left .977 .932 .955 .932 .909 .955 .943 .024
iliac artery right .955 .909 .932 .909 .886 .932 .920 .024
iliac vena left .977 .932 .955 .932 .909 .932 .939 .023
iliac vena right .955 .909 .932 .909 .886 .932 .920 .024
iliopsoas left .943 .887 .925 .925 .906 .925 .918 .019
iliopsoas right .961 .922 .941 .922 .902 .941 .931 .021
inferior vena cava .982 .930 .965 .930 .912 .965 .947 .027
kidney left .906 .887 .906 .849 .868 .943 .893 .033
kidney right .900 .820 .880 .860 .820 .900 .863 .037
liver .945 .909 .964 .909 .891 .945 .927 .028
lung lower lobe left .912 .877 .842 .895 .895 .912 .889 .026
lung lower lobe right .946 .911 .875 .893 .875 .929 .905 .029
lung middle lobe right .939 .918 .878 .959 .837 .980 .918 .053
lung upper lobe left .929 .911 .911 .911 .839 .946 .908 .036
lung upper lobe right .891 .870 .870 .804 .891 .870 .866 .032
pancreas .920 .880 .900 .900 .900 .960 .910 .028
portal and splenic vein .960 .920 .920 .940 .880 .960 .930 .030
pulmonary artery .850 .825 .850 .750 .800 .800 .813 .038
rib left 1 .974 .947 .921 .842 .816 .895 .899 .061
rib left 10 .961 .922 .922 .941 .882 .980 .935 .034
rib left 11 .961 .922 .922 .941 .882 .980 .935 .034
rib left 12 .896 .938 .896 .917 .875 .958 .913 .031
rib left 2 .950 .950 .925 .825 .825 .875 .892 .058
rib left 3 .951 .927 .927 .829 .854 .878 .894 .048
rib left 4 .900 .875 .900 .825 .850 .875 .871 .029
rib left 5 .909 .841 .864 .864 .841 .909 .871 .031
rib left 6 .880 .840 .860 .940 .820 .960 .883 .056
rib left 7 .959 .918 .918 .959 .857 .980 .932 .044
rib left 8 .961 .922 .922 .961 .882 .980 .938 .036
rib left 9 .961 .922 .922 .961 .902 .980 .941 .030
rib right 1 .974 .947 .921 .842 .842 .895 .904 .054
rib right 10 .961 .922 .922 .941 .882 .980 .935 .034
rib right 11 .961 .922 .922 .941 .882 .980 .935 .034
rib right 12 .872 .915 .872 .936 .830 .957 .897 .047
rib right 2 .974 .949 .923 .846 .846 .897 .906 .053
rib right 3 .927 .902 .902 .805 .854 .854 .874 .045
rib right 4 .927 .902 .878 .829 .854 .854 .874 .036
rib right 5 .932 .841 .864 .864 .864 .886 .875 .031
rib right 6 .918 .857 .857 .918 .796 .918 .878 .050
rib right 7 .959 .918 .918 .959 .816 .980 .925 .059
rib right 8 .961 .922 .922 .961 .882 .980 .938 .036
rib right 9 .941 .902 .902 .941 .902 .961 .925 .026
sacrum .955 .909 .955 .909 .909 .932 .928 .022
scapula left .902 .878 .902 .780 .854 .854 .862 .045
scapula right .930 .884 .884 .767 .860 .860 .864 .054
small bowel .938 .854 .896 .917 .896 .896 .899 .028
spleen .980 .940 .940 .940 .900 .980 .947 .030
stomach .980 .941 .941 .941 .902 .980 .948 .030
trachea .951 .927 .902 .805 .805 .854 .874 .063
urinary bladder .977 .953 .953 .907 .907 .930 .938 .028
vertebrae C1 .643 .643 .643 .643 .714 .643 .655 .029
vertebrae C2 .692 .692 .692 .692 .769 .692 .705 .031
vertebrae C3 .643 .714 .571 .714 .857 .714 .702 .095
vertebrae C4 .600 .667 .533 .667 .867 .667 .667 .112
vertebrae C5 .650 .600 .600 .500 .700 .600 .608 .066
vertebrae C6 .818 .758 .788 .758 .606 .636 .727 .086
vertebrae C7 .972 .944 .917 .833 .806 .861 .889 .066
vertebrae L1 .959 .918 .918 .918 .878 .959 .925 .031
vertebrae L2 .909 .886 .909 .886 .886 .977 .909 .035
vertebrae L3 .932 .841 .932 .886 .818 .955 .894 .055
vertebrae L4 .955 .864 .955 .909 .909 .977 .928 .042
vertebrae L5 .953 .884 .953 .907 .907 .953 .926 .031
vertebrae T1 .973 .946 .919 .838 .811 .892 .896 .063
vertebrae T10 .918 .898 .918 .918 .837 .980 .912 .046
vertebrae T11 .958 .917 .917 .938 .875 .979 .931 .036
vertebrae T12 .960 .900 .920 .920 .900 .980 .930 .033
vertebrae T2 .974 .947 .921 .842 .816 .895 .899 .061
vertebrae T3 .947 .921 .895 .816 .816 .868 .877 .054
vertebrae T4 .949 .949 .923 .821 .821 .872 .889 .060
vertebrae T5 .949 .923 .872 .821 .821 .821 .868 .057
vertebrae T6 .944 .944 .917 .833 .833 .889 .894 .051
vertebrae T7 .872 .821 .846 .821 .821 .846 .838 .021
vertebrae T8 .867 .800 .822 .844 .822 .889 .841 .033
vertebrae T9 .878 .857 .857 .898 .796 .939 .871 .048
Average .923 .887 .892 .873 .856 .908
STD .077 .071 .080 .082 .054 .081
3.1.3 Region-based
This section presents the recall of region-based retrieval. An overview of the evaluation is shown in Figure 3. In region-based retrieval, per each anatomical region in the query volume, one volume is retrieved. In the recall computation, the classes present in both the sub-volume of the query and the corresponding retrieved volume are considered TP classes. The classes that are present in the query sub-volume and are missing from the retrieved volume are considered FN.
LABEL:tab:regionbased-recall-29-regions and LABEL:tab:regionbased-recall-104-regions present the retrieval recalls. Compared to volume-based retrieval the average retrieval for the regions is higher. The performance of the models is very close. DreamSim performs slightly better with an average recall of .$7.9$037 for coarse anatomical regions and .$983\pm.032$ for 104 anatomical regions. The retrieval recall for many classes is $1.0$. The standard deviation among classes and the models is low, with the highest standard deviation of .076.076.076.076 and .092.092.092.092, for coarse and fine classes respectively.
Table 6: Region-based retrieval recall of coarse anatomical regions (29 classes) using HNSW Indexing. In each row, bold numbers represent the best-performing values, while italicized numbers indicate the worst-performing. The separate average and standard deviation (STD) columns are color-coded, with blue indicating the best-performing values and yellow indicating the worst-performing values across different models. Additionally, bold numbers in colored columns represent the best classes in terms of average and standard deviation, while italicized values represent the worst-performing class across the models.
Model DINOv1 DINOv2 DreamSim SwinTrans. ResNet50
Dataset (pre-trained) (ImgNet) (ImgNet) (ImgNet) (RadImg) (Fractaldb) (RadImg) Average STD
adrenal gland 1.000 1.000 1.000 .970 .990 1.000 .993 .012
autochthon .992 .992 .992 .992 .992 .992 .992 .000
brain .846 .846 1.000 .923 1.000 1.000 .936 .076
cardiovascular system 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
clavicula .987 .987 1.000 1.000 .961 .987 .987 .014
colon 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
duodenum 1.000 1.000 .979 .958 .958 1.000 .983 .020
esophagus 1.000 1.000 1.000 1.000 1.000 .982 .997 .007
face .882 .882 .824 .824 .882 .824 .853 .032
femur .977 .977 .977 .977 .977 .953 .973 .009
gallbladder .821 .795 .897 .846 .923 .872 .859 .048
gluteus muscles 1.000 1.000 1.000 .992 1.000 .984 .996 .006
hip 1.000 1.000 1.000 1.000 1.000 .989 .998 .005
humerus .931 .931 .977 .966 .897 .977 .946 .032
iliopsoas .980 .990 .990 .980 .980 .990 .985 .005
kidney 1.000 1.000 .980 .971 .941 1.000 .982 .024
liver 1.000 1.000 .982 .982 .945 .982 .982 .020
lung 1.000 1.000 .992 1.000 1.000 1.000 .999 .003
pancreas 1.000 1.000 .980 .980 .980 1.000 .990 .011
portal and splenic vein .980 .980 .980 .980 .980 .980 .980 .000
rib .998 .997 .996 1.000 1.000 1.000 .999 .002
sacrum 1.000 1.000 1.000 .977 1.000 .977 .992 .012
scapula .964 .964 .952 .964 1.000 .988 .972 .018
small bowel .979 .958 .979 .958 .938 .958 .962 .016
spleen 1.000 1.000 .960 1.000 .980 .980 .987 .016
stomach 1.000 1.000 .980 1.000 1.000 1.000 .997 .008
trachea 1.000 1.000 1.000 1.000 1.000 .976 .996 .010
urinary bladder 1.000 1.000 .977 .977 .977 .977 .984 .012
vertebrae 1.000 .999 1.000 .997 .998 .994 .998 .002
Average .977 .976 .979 .973 .976 .978
STD .047 .051 .037 .042 .033 .039
Table 7: Regiond-based retrieval recall of all TS anatomical regions (104 classes) using HNSW Indexing. In each row, bold numbers represent the best-performing values, while italicized numbers indicate the worst-performing. The separate average and standard deviation (STD) columns are color-coded, with blue indicating the best-performing values and yellow indicating the worst-performing values across different models. Additionally, bold numbers in colored columns represent the best classes in terms of average and standard deviation, while italicized values represent the worst-performing class across the models.
Model DINOv1 DINOv2 DreamSim SwinTrans. ResNet50
Dataset (pre-trained) (ImgNet) (ImgNet) (ImgNet) (RadImg) (Fractaldb) (RadImg) Average STD
adrenal gland left .980 1.000 1.000 .940 1.000 1.000 .987 .024
adrenal gland right 1.000 1.000 1.000 .980 .960 1.000 .990 .017
aorta 1.000 1.000 1.000 .984 1.000 1.000 .997 .007
autochthon left 1.000 1.000 1.000 .985 1.000 .985 .995 .008
autochthon right .985 .985 .985 .985 .985 1.000 .987 .006
brain .846 .846 1.000 .923 1.000 1.000 .936 .076
clavicula left .974 .974 1.000 1.000 .974 .974 .983 .013
clavicula right 1.000 1.000 1.000 .974 .947 1.000 .987 .022
colon 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
duodenum 1.000 1.000 .979 .958 .958 1.000 .983 .020
esophagus 1.000 1.000 1.000 1.000 1.000 .982 .997 .007
face .882 .882 .824 .824 .882 .824 .853 .032
femur left .978 .956 .956 .978 .956 .933 .959 .017
femur right .951 1.000 .976 .976 1.000 .976 .980 .018
gallbladder .821 .795 .897 .846 .923 .872 .859 .048
gluteus maximus left 1.000 1.000 1.000 1.000 1.000 .977 .996 .009
gluteus maximus right 1.000 1.000 1.000 1.000 1.000 .977 .996 .009
gluteus medius left 1.000 1.000 1.000 .977 1.000 .977 .992 .012
gluteus medius right 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
gluteus minimus left 1.000 1.000 1.000 .977 1.000 .977 .992 .012
gluteus minimus right 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
heart atrium left 1.000 1.000 1.000 1.000 1.000 .957 .993 .017
heart atrium right 1.000 1.000 1.000 1.000 1.000 .959 .993 .017
heart myocardium 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
heart ventricle left 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
heart ventricle right 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
hip left 1.000 1.000 1.000 1.000 1.000 .977 .996 .009
hip right 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
humerus left .923 .872 .974 .949 .846 .949 .919 .050
humerus right .917 .938 .979 .917 .896 .958 .934 .031
iliac artery left 1.000 1.000 1.000 .977 1.000 .977 .992 .012
iliac artery right 1.000 1.000 1.000 .977 1.000 .977 .992 .012
iliac vena left 1.000 1.000 1.000 .977 1.000 .977 .992 .012
iliac vena right 1.000 1.000 1.000 .977 1.000 .977 .992 .012
iliopsoas left .960 .980 1.000 .980 .980 .980 .980 .013
iliopsoas right .980 .980 .980 .980 .980 1.000 .984 .008
inferior vena cava 1.000 1.000 .982 1.000 .965 1.000 .991 .015
kidney left .981 .943 .981 .962 .943 1.000 .969 .023
kidney right .980 1.000 .980 .980 .939 1.000 .980 .022
liver 1.000 1.000 .982 .982 .945 .982 .982 .020
lung lower lobe left .982 1.000 .982 .982 1.000 .982 .988 .009
lung lower lobe right .982 .982 .982 1.000 1.000 1.000 .991 .010
lung middle lobe right 1.000 1.000 1.000 1.000 .980 .980 .993 .011
lung upper lobe left 1.000 1.000 1.000 1.000 1.000 .982 .997 .007
lung upper lobe right 1.000 .978 .957 .978 1.000 1.000 .986 .018
pancreas 1.000 1.000 .980 .980 .980 1.000 .990 .011
portal and splenic vein .980 .980 .980 .980 .980 .980 .980 .000
pulmonary artery .900 .925 .975 .975 .975 .925 .946 .033
rib left 1 1.000 1.000 1.000 1.000 .974 1.000 .996 .011
rib left 10 .980 .980 1.000 1.000 1.000 1.000 .993 .010
rib left 11 1.000 1.000 .980 .980 1.000 1.000 .993 .010
rib left 12 .938 .896 .979 .979 .958 .979 .955 .033
rib left 2 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
rib left 3 1.000 1.000 1.000 .976 .976 .951 .984 .020
rib left 4 .975 .975 1.000 .950 1.000 1.000 .983 .020
rib left 5 1.000 1.000 1.000 .932 .977 .977 .981 .027
rib left 6 .980 .980 .980 1.000 1.000 .980 .987 .010
rib left 7 1.000 1.000 1.000 1.000 1.000 .980 .997 .008
rib left 8 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
rib left 9 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
rib right 1 1.000 1.000 1.000 1.000 .974 1.000 .996 .011
rib right 10 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
rib right 11 1.000 .980 .980 .980 1.000 1.000 .990 .011
rib right 12 .915 .872 .936 .979 .936 .894 .922 .037
rib right 2 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
rib right 3 .976 .951 .976 .951 .976 .927 .959 .020
rib right 4 .951 .976 1.000 .976 1.000 .976 .980 .018
rib right 5 1.000 .977 1.000 .932 .977 .977 .977 .025
rib right 6 .959 .959 .959 1.000 1.000 .959 .973 .021
rib right 7 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
rib right 8 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
rib right 9 .980 1.000 .980 1.000 1.000 .980 .990 .011
sacrum 1.000 1.000 1.000 .977 1.000 .977 .992 .012
scapula left .976 .976 .951 .976 1.000 1.000 .980 .018
scapula right .953 .953 .953 .953 1.000 .977 .965 .019
small bowel .979 .958 .979 .958 .938 .958 .962 .016
spleen 1.000 1.000 .960 1.000 .980 .980 .987 .016
stomach 1.000 1.000 .980 1.000 1.000 1.000 .997 .008
trachea 1.000 1.000 1.000 1.000 1.000 .976 .996 .010
urinary bladder 1.000 1.000 .977 .977 .977 .977 .984 .012
vertebrae C1 .929 .857 .929 .929 .929 .857 .905 .037
vertebrae C2 1.000 .923 1.000 1.000 .846 .923 .949 .063
vertebrae C3 .929 .857 1.000 1.000 .929 1.000 .952 .058
vertebrae C4 .867 .800 .933 .933 .733 .733 .833 .092
vertebrae C5 .850 .750 .850 .850 .900 .900 .850 .055
vertebrae C6 .909 .848 .848 .939 .788 .848 .864 .053
vertebrae C7 1.000 1.000 1.000 .972 .861 1.000 .972 .056
vertebrae L1 1.000 1.000 .938 1.000 1.000 1.000 .990 .026
vertebrae L2 .955 1.000 .977 .932 .955 .977 .966 .024
vertebrae L3 .977 1.000 .977 .977 .909 1.000 .973 .033
vertebrae L4 1.000 .932 1.000 1.000 .977 1.000 .985 .028
vertebrae L5 1.000 .953 1.000 1.000 .953 1.000 .984 .024
vertebrae T1 1.000 1.000 1.000 .973 .919 1.000 .982 .033
vertebrae T10 .980 1.000 1.000 .980 1.000 .980 .990 .011
vertebrae T11 .979 1.000 1.000 .979 .979 1.000 .990 .011
vertebrae T12 1.000 .980 1.000 1.000 .980 1.000 .993 .011
vertebrae T2 1.000 1.000 1.000 1.000 .974 1.000 .996 .011
vertebrae T3 .974 .974 .974 .974 .974 .974 .974 .000
vertebrae T4 1.000 1.000 1.000 1.000 1.000 .974 .996 .010
vertebrae T5 .974 .974 .974 1.000 1.000 .949 .979 .019
vertebrae T6 .944 .944 1.000 1.000 1.000 .944 .972 .030
vertebrae T7 .974 .947 .947 .974 1.000 .947 .965 .021
vertebrae T8 .978 .978 .956 .956 .956 .933 .959 .017
vertebrae T9 1.000 .959 .980 .980 1.000 .959 .980 .018
Average .979 .972 .983 .978 .973 .974
STD .037 .050 .032 .032 .046 .042
3.1.4 Localized
This section presents the recall and localization-ratio of localized retrieval. An overview of the evaluation is shown in Figure 4. In localized retrieval, per each anatomical region in the query volume, one volume is retrieved.
Localized Retrieval Recall
The recall calculation for localized retrieval is explained in Section 2.4.3 and an overview is shown in Figure 4. Table 8 and LABEL:tab:localized-recall-104-regions present the retrieval recalls. Compared to region-based retrieval the average retrieval for regions is lower which is expected based on the more strict metric defined. The performance of models is close, especially the self-supervised models. DINOv2 performs best for 29 anatomical regions with an average recall of .$941\pm.077$. For 104 regions the performance of models is even closer with DINOv1 performing slightly better with an average recall of .$929\pm.085$.
Table 8: Localized retrieval recall of coarse anatomical regions (29 classes) using HNSW Indexing, $L=15$. In each row, bold numbers represent the best-performing values, while italicized numbers indicate the worst-performing. The separate average and standard deviation (STD) columns are color-coded, with blue indicating the best-performing values and yellow indicating the worst-performing values across different models. Additionally, bold numbers in colored columns represent the best classes in terms of average and standard deviation, while italicized values represent the worst-performing class across the models.
Model DINOv1 DINOv2 DreamSim SwinTrans. ResNet50
Dataset (pre-trained) (ImgNet) (ImgNet) (ImgNet) (RadImg) (Fractaldb) (RadImg) Average STD
adrenal gland .960 .920 .920 .850 .790 .860 .883 .062
autochthon .985 .992 .969 .985 .954 .992 .979 .015
brain .846 .846 1.000 .923 .846 1.000 .910 .076
cardiovascular system 1.000 .998 .995 .997 .983 .976 .991 .010
clavicula .935 .987 .974 .961 .857 .987 .950 .050
colon .962 1.000 1.000 .942 .923 .962 .965 .031
duodenum .875 .958 .917 .854 .813 .917 .889 .052
esophagus .964 .964 .982 1.000 .964 .946 .970 .018
face .765 .824 .824 .824 .882 .824 .824 .037
femur .860 .977 .965 .965 .977 .942 .948 .045
gallbladder .513 .641 .538 .538 .538 .513 .547 .048
gluteus muscles 1.000 .988 1.000 .988 .977 .957 .985 .016
hip 1.000 1.000 1.000 .977 .989 .966 .989 .014
humerus .747 .793 .862 .701 .586 .736 .738 .093
iliopsoas .980 .950 .970 .950 .970 .941 .960 .015
kidney .971 .951 .922 .922 .863 .941 .928 .037
liver .982 .964 .909 .945 .891 .927 .936 .034
lung 1.000 1.000 .973 .992 .962 .943 .979 .023
pancreas .918 .959 .898 .918 .755 .918 .895 .071
portal and splenic vein .940 .960 .920 .860 .780 .760 .870 .085
rib .998 .995 .987 .990 .971 .981 .987 .010
sacrum .955 .955 1.000 .977 .955 .932 .962 .023
scapula .952 .964 .929 .952 .929 .893 .937 .026
small bowel .979 .938 .938 .958 .896 .938 .941 .028
spleen .980 .920 .900 .960 .900 .880 .923 .039
stomach .980 .961 .941 .941 .941 .922 .948 .020
trachea 1.000 .951 .976 1.000 .927 .927 .963 .034
urinary bladder .977 .930 .930 .884 .860 .860 .907 .047
vertebrae .999 .995 .993 .990 .971 .981 .988 .010
Average .932 .941 .936 .922 .884 .908
STD .105 .077 .089 .099 .111 .099
Table 9: Localized retrieval recall of all TS anatomical regions (104 classes) using HNSW Indexing, $L=15$. In each row, bold numbers represent the best-performing values, while italicized numbers indicate the worst-performing. The separate average and standard deviation (STD) columns are color-coded, with blue indicating the best-performing values and yellow indicating the worst-performing values across different models. Additionally, bold numbers in colored columns represent the best classes in terms of average and standard deviation, while italicized values represent the worst-performing class across the models.
Model DINOv1 DINOv2 DreamSim SwinTrans. ResNet50
Dataset (pre-trained) (ImgNet) (ImgNet) (ImgNet) (RadImg) (Fractaldb) (RadImg) Average STD
adrenal gland left .900 .920 .860 .740 .760 .740 .820 .083
adrenal gland right .880 .840 .880 .840 .620 .840 .817 .098
aorta 1.000 .967 1.000 .967 .984 .967 .981 .016
autochthon left .985 1.000 .969 .985 .954 .985 .979 .016
autochthon right .985 .985 .969 .985 .954 1.000 .979 .016
brain .846 .846 1.000 .923 .846 1.000 .910 .076
clavicula left .923 .974 .949 .949 .872 .949 .936 .035
clavicula right .921 1.000 1.000 .947 .842 1.000 .952 .063
colon .962 1.000 1.000 .942 .923 .962 .965 .031
duodenum .875 .958 .917 .854 .813 .917 .889 .052
esophagus .964 .964 .982 1.000 .964 .946 .970 .018
face .765 .824 .824 .824 .882 .824 .824 .037
femur left .844 .956 .933 .956 .956 .911 .926 .044
femur right .878 1.000 .976 .976 1.000 .976 .967 .045
gallbladder .513 .641 .538 .538 .538 .513 .547 .048
gluteus maximus left 1.000 1.000 1.000 .977 .930 .953 .977 .029
gluteus maximus right 1.000 1.000 .977 .977 .977 .953 .981 .018
gluteus medius left 1.000 .977 .977 .977 .932 .932 .966 .028
gluteus medius right .977 .977 1.000 .953 1.000 .953 .977 .021
gluteus minimus left .977 .977 1.000 .977 .907 .953 .965 .032
gluteus minimus right 1.000 .952 .976 1.000 .976 1.000 .984 .019
heart atrium left .872 .851 .830 .936 .851 .638 .830 .101
heart atrium right .980 .918 .980 .939 .837 .918 .929 .053
heart myocardium 1.000 .980 1.000 .959 .898 .898 .956 .047
heart ventricle left .980 .939 .959 .939 .898 .918 .939 .029
heart ventricle right .980 1.000 .980 .959 .918 .878 .952 .046
hip left 1.000 1.000 1.000 .977 .977 .955 .985 .019
hip right 1.000 1.000 1.000 .955 1.000 .977 .989 .019
humerus left .744 .744 .846 .692 .538 .744 .718 .101
humerus right .750 .708 .833 .604 .542 .667 .684 .104
iliac artery left .977 1.000 1.000 .955 .932 .932 .966 .031
iliac artery right 1.000 1.000 1.000 .955 .955 .932 .973 .030
iliac vena left .977 .955 1.000 .977 .955 .955 .970 .019
iliac vena right .977 .955 .977 .955 .977 .932 .962 .019
iliopsoas left .960 .960 .980 .940 .960 .920 .953 .021
iliopsoas right .980 .941 .961 .941 .961 .961 .958 .015
inferior vena cava 1.000 .965 .965 .982 .947 .982 .974 .018
kidney left .943 .868 .906 .868 .830 .906 .887 .040
kidney right .918 .959 .918 .918 .816 .878 .901 .049
liver .982 .964 .909 .945 .891 .927 .936 .034
lung lower lobe left .965 .930 .965 .965 .930 .895 .942 .029
lung lower lobe right .982 .929 .911 .911 .946 .893 .929 .032
lung middle lobe right .980 1.000 .959 .959 .878 .898 .946 .048
lung upper lobe left .982 .982 .982 1.000 .893 .893 .955 .049
lung upper lobe right .935 .957 .913 .935 .870 .848 .909 .042
pancreas .918 .959 .898 .918 .755 .918 .895 .071
portal vein and splenic vein .940 .960 .920 .860 .780 .760 .870 .085
pulmonary artery .825 .850 .925 .825 .725 .775 .821 .068
rib left 1 .974 .974 1.000 .974 .921 .974 .969 .026
rib left 10 .980 .961 .961 .941 .902 .922 .944 .029
rib left 11 .980 1.000 .961 .902 .980 .980 .967 .034
rib left 12 .833 .813 .854 .875 .771 .875 .837 .040
rib left 2 .975 .975 .950 .975 .850 .850 .929 .062
rib left 3 .951 .927 .902 .927 .951 .829 .915 .046
rib left 4 .950 .925 .950 .950 .950 .925 .942 .013
rib left 5 .886 .864 .886 .909 .864 .864 .879 .019
rib left 6 .960 .900 .860 .940 .860 .880 .900 .042
rib left 7 1.000 .980 .959 .959 .898 .939 .956 .035
rib left 8 1.000 1.000 .902 .961 .804 .961 .938 .075
rib left 9 .980 1.000 .941 .941 .882 .941 .948 .041
rib right 1 1.000 .947 1.000 .974 .895 .947 .961 .040
rib right 10 .980 1.000 1.000 .922 .922 .961 .964 .036
rib right 11 1.000 .980 .980 .882 .961 .980 .964 .042
rib right 12 .766 .787 .809 .830 .809 .809 .801 .022
rib right 2 1.000 .974 .949 .974 .897 .949 .957 .035
rib right 3 .951 .927 .902 .902 .927 .805 .902 .051
rib right 4 .951 .976 .927 .927 .951 .854 .931 .042
rib right 5 .864 .886 .886 .864 .864 .841 .867 .017
rib right 6 .939 .918 .857 .918 .939 .857 .905 .038
rib right 7 1.000 1.000 .918 .959 .898 .959 .956 .042
rib right 8 1.000 .941 .980 .980 .882 .961 .958 .042
rib right 9 .980 1.000 .941 .961 .882 .922 .948 .042
sacrum .955 .955 1.000 .977 .955 .932 .962 .023
scapula left .951 .951 .927 .976 .927 .902 .939 .026
scapula right .953 .953 .930 .930 .930 .884 .930 .025
small bowel .979 .938 .938 .958 .896 .938 .941 .028
spleen .980 .920 .900 .960 .900 .880 .923 .039
stomach .980 .961 .941 .941 .941 .922 .948 .020
trachea 1.000 .951 .976 1.000 .927 .927 .963 .034
urinary bladder .977 .930 .930 .884 .860 .860 .907 .047
vertebrae C1 .929 .786 .857 .857 .643 .643 .786 .120
vertebrae C2 .923 .923 1.000 .846 .846 .846 .897 .063
vertebrae C3 .857 .714 .786 .714 .643 .500 .702 .123
vertebrae C4 .667 .600 .600 .733 .267 .467 .556 .167
vertebrae C5 .800 .550 .800 .750 .400 .450 .625 .181
vertebrae C6 .667 .636 .727 .515 .242 .606 .566 .173
vertebrae C7 .861 .861 .833 .750 .500 .694 .750 .139
vertebrae L1 .896 .896 .813 .854 .771 .813 .840 .050
vertebrae L2 .795 .932 .932 .841 .682 .886 .845 .096
vertebrae L3 .886 .818 .841 .886 .727 .932 .848 .071
vertebrae L4 .977 .773 .886 .977 .795 .886 .883 .087
vertebrae L5 .977 .907 1.000 .930 .884 .907 .934 .045
vertebrae T1 .973 .973 .973 .919 .865 .892 .932 .048
vertebrae T10 .878 .878 .898 .898 .673 .796 .837 .088
vertebrae T11 .833 .833 .813 .750 .729 .708 .778 .055
vertebrae T12 .939 .918 .878 .898 .694 .796 .854 .093
vertebrae T2 .974 .921 .974 1.000 .763 .895 .921 .086
vertebrae T3 .921 .921 .947 .895 .868 .789 .890 .056
vertebrae T4 1.000 .923 .872 .872 .795 .821 .880 .074
vertebrae T5 .974 .923 .974 .846 .744 .718 .863 .113
vertebrae T6 .944 .861 .972 .861 .778 .778 .866 .081
vertebrae T7 .895 .921 .921 .921 .763 .711 .855 .094
vertebrae T8 .778 .822 .889 .889 .667 .711 .793 .092
vertebrae T9 .878 .857 .898 .796 .735 .735 .816 .072
Average .929 .917 .924 .906 .840 .866
STD .085 .091 .078 .092 .145 .118
Localization-ratio
The localization-ratio is computed based on (2). This measure shows how many slices that contributed to the retrieval of the volume actually contained the desired organ. Table 10 and LABEL:tab:localization-ratio-104-regions show the localization-ratio for 29 coarse and 104 TS original classes. DreamSim shows the best average localization-ratio with an average localization-ratio of .$864\pm.145$ and .$803\pm.130$ for coarse and original TS classes, respectively.
Table 10: Localization-ratio of coarse anatomical regions (29 classes) using HNSW Indexing, $L=15$. In each row, bold numbers represent the best-performing values, while italicized numbers indicate the worst-performing. The separate average and standard deviation (STD) columns are color-coded, with blue indicating the best-performing values and yellow indicating the worst-performing values across different models. Additionally, bold numbers in colored columns represent the best classes in terms of average and standard deviation, while italicized values represent the worst-performing class across the models.
Model DINOv1 DINOv2 DreamSim SwinTrans. ResNet50
Dataset (pre-trained) (ImgNet) (ImgNet) (ImgNet) (RadImg) (Fractaldb) (RadImg) Average STD
adrenal gland .801 .651 .724 .669 .525 .617 .665 .094
autochthon .971 .968 .960 .973 .938 .963 .962 .013
brain .842 .837 .941 .869 .835 .873 .866 .040
cardiovascular system .998 .993 .995 .995 .954 .954 .982 .021
clavicula .890 .945 .961 .923 .788 .952 .910 .065
colon .914 .963 .964 .865 .831 .892 .905 .053
duodenum .627 .696 .765 .656 .654 .747 .691 .055
esophagus .895 .913 .931 .961 .924 .858 .914 .035
face .748 .774 .787 .733 .804 .736 .763 .029
femur .873 .948 .990 .948 .925 .960 .940 .039
gallbladder .219 .388 .302 .320 .363 .369 .327 .062
gluteus muscles .986 .954 .990 .991 .976 .989 .981 .014
hip .978 .951 .976 .945 .951 .931 .955 .018
humerus .592 .581 .749 .583 .452 .571 .588 .094
iliopsoas .929 .912 .937 .912 .909 .900 .916 .014
kidney .813 .819 .804 .859 .757 .824 .813 .033
liver .873 .841 .830 .831 .832 .860 .844 .018
lung .979 .976 .962 .969 .928 .878 .949 .039
pancreas .697 .787 .778 .708 .573 .770 .719 .081
portal and splenic vein .725 .652 .692 .630 .552 .493 .624 .087
rib .964 .970 .999 .972 .972 .933 .968 .021
sacrum .875 .860 .962 .889 .932 .855 .895 .043
scapula .936 .897 .902 .889 .908 .821 .892 .038
small bowel .905 .818 .889 .881 .847 .895 .872 .033
spleen .864 .742 .736 .735 .815 .710 .767 .059
stomach .876 .780 .827 .815 .833 .807 .823 .032
trachea .907 .874 .915 .912 .797 .805 .868 .055
urinary bladder .752 .709 .804 .776 .670 .685 .733 .053
vertebrae .998 .998 .972 .998 .970 .995 .988 .014
Average .842 .834 .864 .835 .800 .815
STD .161 .144 .145 .155 .168 .152
Table 11: Localization-ratio of all TS anatomical regions (104 classes) using HNSW Indexing, $L=15$. In each row, bold numbers represent the best-performing values, while italicized numbers indicate the worst-performing. The separate average and standard deviation (STD) columns are color-coded, with blue indicating the best-performing values and yellow indicating the worst-performing values across different models. Additionally, bold numbers in colored columns represent the best classes in terms of average and standard deviation, while italicized values represent the worst-performing class across the models.
Model DINOv1 DINOv2 DreamSim SwinTrans. ResNet50
Dataset (pre-trained) (ImgNet) (ImgNet) (ImgNet) (RadImg) (Fractaldb) (RadImg) Average STD
adrenal gland left .683 .569 .602 .526 .488 .431 .550 .089
adrenal gland right .654 .525 .641 .578 .410 .535 .557 .089
aorta .943 .921 .967 .936 .927 .889 .931 .026
autochthon left .969 .980 .960 .983 .938 .964 .966 .016
autochthon right .972 .956 .954 .968 .938 .964 .959 .012
brain .842 .837 .941 .869 .835 .873 .866 .040
clavicula left .871 .908 .891 .882 .827 .885 .877 .027
clavicula right .859 .924 .923 .886 .752 .924 .878 .067
colon .914 .963 .964 .865 .831 .892 .905 .053
duodenum .627 .696 .765 .656 .654 .747 .691 .055
esophagus .895 .913 .931 .961 .924 .858 .914 .035
face .748 .774 .787 .733 .804 .736 .763 .029
femur left .836 .907 .908 .928 .868 .862 .885 .035
femur right .870 .922 .949 .951 .916 .947 .926 .031
gallbladder .219 .388 .302 .320 .363 .369 .327 .062
gluteus maximus left .981 .921 .956 .947 .868 .896 .928 .041
gluteus maximus right .992 .928 .945 .916 .937 .913 .938 .029
gluteus medius left .966 .877 .947 .936 .904 .865 .916 .040
gluteus medius right .932 .911 .973 .912 .944 .900 .929 .027
gluteus minimus left .907 .834 .907 .903 .834 .899 .881 .036
gluteus minimus right .915 .790 .852 .906 .861 .947 .878 .056
heart atrium left .727 .703 .756 .671 .672 .452 .664 .109
heart atrium right .866 .785 .912 .794 .787 .794 .823 .053
heart myocardium .881 .794 .903 .840 .815 .782 .836 .048
heart ventricle left .850 .764 .879 .864 .807 .814 .830 .043
heart ventricle right .857 .833 .876 .839 .850 .769 .837 .037
hip left .981 .949 .974 .929 .952 .913 .950 .026
hip right .980 .941 .972 .941 .972 .948 .959 .017
humerus left .559 .619 .679 .625 .438 .583 .584 .082
humerus right .567 .563 .686 .515 .390 .542 .544 .096
iliac artery left .888 .895 .918 .933 .890 .847 .895 .030
iliac artery right .929 .885 .935 .911 .916 .833 .901 .038
iliac vena left .906 .869 .927 .916 .867 .860 .891 .029
iliac vena right .918 .883 .940 .901 .904 .839 .897 .034
iliopsoas left .920 .926 .954 .890 .871 .870 .905 .034
iliopsoas right .922 .908 .934 .888 .885 .900 .906 .019
inferior vena cava .944 .927 .894 .936 .922 .912 .923 .018
kidney left .764 .663 .798 .717 .637 .763 .724 .063
kidney right .720 .751 .773 .752 .690 .783 .745 .035
liver .873 .841 .830 .831 .832 .860 .844 .018
lung lower lobe left .876 .879 .873 .865 .835 .813 .857 .027
lung lower lobe right .912 .855 .866 .865 .817 .779 .849 .046
lung middle lobe right .802 .832 .830 .856 .750 .738 .802 .048
lung upper lobe left .921 .938 .925 .945 .831 .844 .901 .050
lung upper lobe right .887 .879 .881 .873 .848 .779 .858 .041
pancreas .697 .787 .778 .708 .573 .770 .719 .081
portal and splenic vein .725 .652 .692 .630 .552 .493 .624 .087
pulmonary artery .689 .697 .758 .647 .493 .575 .643 .095
rib left 1 .861 .878 .922 .848 .827 .826 .860 .036
rib left 10 .878 .792 .823 .829 .815 .790 .821 .032
rib left 11 .820 .768 .827 .799 .854 .861 .821 .034
rib left 12 .656 .480 .697 .643 .561 .649 .614 .080
rib left 2 .860 .801 .819 .785 .736 .770 .795 .043
rib left 3 .837 .806 .799 .824 .854 .743 .811 .038
rib left 4 .801 .829 .861 .816 .820 .816 .824 .020
rib left 5 .785 .763 .810 .774 .732 .753 .769 .027
rib left 6 .824 .789 .768 .774 .723 .789 .778 .033
rib left 7 .887 .888 .842 .856 .771 .808 .842 .046
rib left 8 .847 .881 .826 .826 .717 .819 .819 .055
rib left 9 .831 .862 .822 .860 .758 .845 .830 .039
rib right 1 .864 .826 .871 .790 .826 .857 .839 .031
rib right 10 .890 .810 .826 .820 .825 .840 .835 .029
rib right 11 .863 .785 .855 .792 .831 .794 .820 .034
rib right 12 .554 .488 .643 .677 .536 .562 .577 .070
rib right 2 .883 .817 .834 .844 .726 .850 .826 .053
rib right 3 .867 .814 .803 .827 .775 .724 .801 .049
rib right 4 .850 .836 .880 .813 .842 .766 .831 .038
rib right 5 .750 .774 .819 .753 .761 .746 .767 .027
rib right 6 .793 .790 .789 .757 .773 .729 .772 .025
rib right 7 .849 .878 .810 .853 .754 .824 .828 .043
rib right 8 .867 .840 .868 .846 .763 .834 .836 .039
rib right 9 .860 .855 .804 .865 .790 .814 .831 .033
sacrum .875 .860 .962 .889 .932 .855 .895 .043
scapula left .891 .869 .869 .909 .892 .850 .880 .021
scapula right .936 .875 .907 .896 .895 .823 .889 .038
small bowel .905 .818 .889 .881 .847 .895 .872 .033
spleen .864 .742 .736 .735 .815 .710 .767 .059
stomach .876 .780 .827 .815 .833 .807 .823 .032
trachea .907 .874 .915 .912 .797 .805 .868 .055
urinary bladder .752 .709 .804 .776 .670 .685 .733 .053
vertebrae C1 .581 .582 .621 .509 .502 .564 .560 .046
vertebrae C2 .713 .719 .828 .595 .656 .576 .681 .093
vertebrae C3 .670 .659 .597 .524 .433 .243 .521 .163
vertebrae C4 .435 .350 .475 .497 .267 .276 .383 .100
vertebrae C5 .559 .408 .613 .631 .330 .286 .471 .149
vertebrae C6 .525 .495 .504 .360 .183 .555 .437 .141
vertebrae C7 .719 .679 .704 .618 .382 .623 .621 .124
vertebrae L1 .645 .579 .613 .655 .536 .582 .602 .045
vertebrae L2 .542 .599 .659 .579 .535 .700 .602 .066
vertebrae L3 .720 .494 .579 .593 .595 .655 .606 .076
vertebrae L4 .638 .381 .619 .721 .637 .642 .606 .116
vertebrae L5 .782 .651 .757 .773 .722 .698 .731 .050
vertebrae T1 .741 .745 .634 .700 .663 .731 .702 .046
vertebrae T10 .553 .545 .613 .649 .489 .556 .568 .056
vertebrae T11 .580 .528 .538 .470 .447 .498 .510 .049
vertebrae T12 .564 .553 .624 .638 .490 .538 .568 .055
vertebrae T2 .814 .724 .758 .742 .669 .728 .739 .047
vertebrae T3 .794 .732 .792 .740 .683 .598 .723 .074
vertebrae T4 .739 .728 .700 .713 .561 .622 .677 .070
vertebrae T5 .684 .687 .775 .700 .536 .497 .647 .107
vertebrae T6 .674 .596 .595 .585 .536 .525 .585 .053
vertebrae T7 .706 .743 .686 .583 .507 .543 .628 .097
vertebrae T8 .572 .538 .756 .588 .432 .438 .554 .119
vertebrae T9 .633 .589 .755 .550 .578 .522 .605 .083
Average .793 .763 .803 .773 .722 .736
STD .140 .148 .130 .143 .178 .162
3.2 Re-ranking
This section presents the retrieval recalls after applying the re-ranking method of Section 2.5.
3.2.1 Volume-based
LABEL:tab:volumewise-recall-29-regions-reranking and LABEL:tab:volumewise-recall-104-regions-reranking show the retrieval recalls for 29 coarse anatomical structures and 104 original TS anatomical structures using the proposed re-ranking method. All the recalls are improved using re-ranking. The performance of the models for 29 classes is close with only slight differences. DINOv1 and DreamSim have a slightly better recall in comparison, with an average recall of .967.967.967.967 but the standard deviation of DINOv1 is slightly lower (.040.040.040.040 vs. .045.045.045.045). In 104 anatomical regions, SwinTransformer performs better than the other models with an average recall of .924.924.924.924 but its standard deviation (.072.072.072.072) is the lowest.
Table 12: Volume-based retrieval recall of coarse anatomical regions (29 classes) using HNSW Indexing and re-ranking. In each row, bold numbers represent the best-performing values, while italicized numbers indicate the worst-performing. The separate average and standard deviation (STD) columns are color-coded, with blue indicating the best-performing values and yellow indicating the worst-performing values across different models. Additionally, bold numbers in colored columns represent the best classes in terms of average and standard deviation, while italicized values represent the worst-performing class across the models.
Model DINOv1 DINOv2 DreamSim SwinTrans. ResNet50
Dataset (pre-trained) (ImgNet) (ImgNet) (ImgNet) (RadImg) (Fractaldb) (RadImg) Average STD
adrenal gland 1.000 1.000 .980 .960 1.000 .960 .983 .020
autochthon .985 .969 .985 .985 1.000 1.000 .987 .012
brain .923 .923 .846 .923 .692 .846 .859 .090
cardiovascular system 1.000 1.000 .984 1.000 1.000 1.000 .997 .006
clavicula .974 .974 .974 .974 1.000 1.000 .983 .013
colon 1.000 .981 1.000 .962 .962 .981 .981 .017
duodenum .920 .900 .960 .940 1.000 .920 .940 .036
esophagus .982 1.000 .982 1.000 1.000 1.000 .994 .009
face .941 .882 .824 .765 .647 .824 .814 .092
femur .956 .978 .933 .956 .978 .933 .956 .020
gallbladder .821 .821 .897 .872 .872 .821 .850 .034
gluteus muscles 1.000 1.000 1.000 1.000 1.000 .977 .996 .009
hip 1.000 1.000 1.000 1.000 1.000 .977 .996 .009
humerus .918 .857 .959 .918 .918 .980 .925 .042
iliopsoas .962 .962 1.000 1.000 1.000 .943 .978 .025
kidney .964 .945 1.000 .964 .982 .964 .970 .019
liver .982 .964 1.000 .982 .982 1.000 .985 .014
lung .983 .983 .948 .966 .983 .983 .974 .014
pancreas .940 .960 .980 .980 .980 .960 .967 .016
portal and splenic vein 1.000 .980 .980 .980 .980 .980 .983 .008
rib .983 .983 .983 .983 .983 1.000 .986 .007
sacrum .977 .977 1.000 .977 1.000 .977 .985 .012
scapula .909 .932 .932 .909 .977 .955 .936 .027
small bowel .958 .958 .958 .958 1.000 .938 .962 .020
spleen 1.000 1.000 .980 .980 1.000 1.000 .993 .010
stomach 1.000 1.000 1.000 .980 1.000 .980 .993 .010
trachea .976 .976 .951 .976 .951 .976 .967 .013
urinary bladder 1.000 1.000 1.000 1.000 1.000 .977 .996 .009
vertebrae .984 .969 1.000 .984 1.000 1.000 .990 .013
Average .967 .961 .967 .961 .962 .960
STD .040 .045 .045 .049 .086 .050
Table 13: Volume-based retrieval recall of all TS anatomical regions (104 classes) using HNSW Indexing and re-ranking. In each row, bold numbers represent the best-performing values, while italicized numbers indicate the worst-performing. The separate average and standard deviation (STD) columns are color-coded, with blue indicating the best-performing values and yellow indicating the worst-performing values across different models. Additionally, bold numbers in colored columns represent the best classes in terms of average and standard deviation, while italicized values represent the worst-performing class across the models.
Model DINOv1 DINOv2 DreamSim SwinTrans. ResNet50
Dataset (pre-trained) (ImgNet) (ImgNet) (ImgNet) (RadImg) (Fractaldb) (RadImg) Average STD
adrenal gland left .920 .880 .880 .940 .920 .880 .903 .027
adrenal gland right .940 .920 .880 .940 .920 .900 .917 .023
aorta .951 .902 .885 .951 .885 .918 .915 .030
autochthon left .938 .877 .892 .938 .908 .923 .913 .025
autochthon right .938 .877 .892 .938 .908 .923 .913 .025
brain .923 .846 .846 .769 .692 .769 .808 .081
clavicula left .949 .897 .949 .897 .974 .974 .940 .035
clavicula right .974 .921 .974 .921 .974 .974 .956 .027
colon .943 .887 .887 .943 .868 .906 .906 .032
duodenum .860 .800 .840 .920 .880 .840 .857 .041
esophagus .929 .911 .911 .946 .929 .929 .926 .013
face .941 .824 .824 .647 .647 .765 .775 .114
femur left .889 .889 .800 .933 .867 .844 .870 .045
femur right .878 .878 .805 .976 .902 .854 .882 .057
gallbladder .795 .744 .795 .872 .821 .769 .799 .044
gluteus maximus left .930 .907 .860 .977 .907 .884 .911 .040
gluteus maximus right .930 .907 .837 .977 .907 .884 .907 .047
gluteus medius left .932 .886 .864 .977 .909 .886 .909 .041
gluteus medius right .930 .884 .860 .977 .907 .907 .911 .040
gluteus minimus left .930 .907 .860 .977 .907 .884 .911 .040
gluteus minimus right .929 .905 .857 .976 .905 .905 .913 .039
heart atrium left .894 .915 .915 .979 .936 .936 .929 .029
heart atrium right .918 .918 .898 .980 .918 .939 .929 .028
heart myocardium .918 .918 .898 .980 .918 .939 .929 .028
heart ventricle left .918 .918 .898 .980 .918 .939 .929 .028
heart ventricle right .918 .918 .898 .980 .918 .939 .929 .028
hip left .932 .886 .864 .977 .909 .886 .909 .041
hip right .932 .886 .864 .977 .909 .886 .909 .041
humerus left .949 .821 .949 .923 .974 .949 .927 .055
humerus right .854 .771 .875 .875 .854 .896 .854 .044
iliac artery left .932 .886 .864 .977 .909 .886 .909 .041
iliac artery right .909 .864 .841 .955 .909 .886 .894 .040
iliac vena left .932 .886 .864 .977 .909 .886 .909 .041
iliac vena right .909 .864 .841 .955 .909 .886 .894 .040
iliopsoas left .906 .868 .887 .981 .887 .868 .899 .042
iliopsoas right .902 .843 .882 .980 .882 .882 .895 .046
inferior vena cava .947 .895 .895 .965 .877 .930 .918 .034
kidney left .868 .849 .868 .925 .868 .887 .877 .026
kidney right .860 .800 .880 .940 .860 .860 .867 .045
liver .927 .891 .891 .964 .873 .927 .912 .033
lung lower lobe left .912 .877 .842 .895 .912 .895 .889 .026
lung lower lobe right .929 .875 .857 .911 .911 .911 .899 .027
lung middle lobe right .918 .918 .898 .980 .918 .939 .929 .028
lung upper lobe left .929 .911 .911 .929 .929 .929 .923 .009
lung upper lobe right .891 .848 .870 .870 .935 .913 .888 .032
pancreas .880 .880 .880 .960 .880 .900 .897 .032
portal and splenic vein .940 .900 .880 .960 .900 .920 .917 .029
pulmonary artery .850 .850 .825 .850 .875 .875 .854 .019
rib left 1 .974 .921 .974 .921 .974 .974 .956 .027
rib left 10 .922 .902 .882 .961 .902 .941 .918 .029
rib left 11 .922 .902 .882 .961 .902 .941 .918 .029
rib left 12 .896 .917 .875 .917 .896 .896 .899 .016
rib left 2 .950 .925 .950 .925 .950 .950 .942 .013
rib left 3 .951 .902 .951 .927 .951 .951 .939 .020
rib left 4 .925 .900 .900 .900 .950 .925 .917 .020
rib left 5 .886 .886 .886 .909 .977 .932 .913 .036
rib left 6 .860 .860 .880 .960 .900 .900 .893 .037
rib left 7 .939 .918 .918 .980 .918 .939 .935 .024
rib left 8 .941 .922 .902 .961 .922 .941 .931 .021
rib left 9 .941 .922 .902 .961 .922 .941 .931 .021
rib right 1 .974 .921 .974 .921 .974 .974 .956 .027
rib right 10 .922 .902 .882 .961 .902 .941 .918 .029
rib right 11 .922 .902 .882 .961 .902 .941 .918 .029
rib right 12 .851 .894 .851 .957 .872 .894 .887 .040
rib right 2 .974 .923 .974 .923 .974 .974 .957 .026
rib right 3 .927 .878 .927 .902 .951 .927 .919 .025
rib right 4 .927 .878 .878 .902 .927 .927 .907 .024
rib right 5 .886 .864 .886 .932 .977 .932 .913 .042
rib right 6 .878 .837 .898 .959 .857 .898 .888 .042
rib right 7 .939 .918 .918 .980 .918 .939 .935 .024
rib right 8 .941 .922 .902 .961 .922 .941 .931 .021
rib right 9 .922 .902 .882 .941 .922 .922 .915 .020
sacrum .909 .864 .864 .955 .909 .886 .898 .034
scapula left .902 .902 .902 .878 .951 .902 .907 .024
scapula right .907 .860 .884 .860 .953 .930 .899 .038
small bowel .896 .854 .833 .938 .896 .854 .878 .038
spleen .940 .920 .880 .960 .920 .940 .927 .027
stomach .941 .922 .902 .961 .902 .922 .925 .023
trachea .951 .902 .927 .902 .927 .927 .923 .018
urinary bladder .930 .907 .860 .977 .907 .884 .911 .040
vertebrae C1 .857 .786 .857 .714 .714 .786 .786 .064
vertebrae C2 .923 .846 .923 .769 .769 .846 .846 .069
vertebrae C3 .929 .857 .929 .714 .857 .786 .845 .084
vertebrae C4 .867 .800 .867 .667 .800 .733 .789 .078
vertebrae C5 .750 .700 .750 .600 .650 .650 .683 .061
vertebrae C6 .788 .848 .909 .788 .758 .758 .808 .060
vertebrae C7 .972 .917 .972 .917 1.000 .944 .954 .034
vertebrae L1 .898 .878 .878 .939 .898 .918 .901 .024
vertebrae L2 .864 .841 .886 .955 .886 .841 .879 .042
vertebrae L3 .886 .818 .864 .932 .864 .841 .867 .039
vertebrae L4 .909 .841 .864 .932 .886 .886 .886 .032
vertebrae L5 .907 .837 .860 .953 .907 .907 .895 .041
vertebrae T1 .973 .919 .973 .919 .973 .973 .955 .028
vertebrae T10 .898 .918 .898 .959 .918 .939 .922 .024
vertebrae T11 .938 .917 .917 .979 .917 .938 .934 .024
vertebrae T12 .940 .920 .900 .940 .920 .940 .927 .016
vertebrae T2 .974 .921 .974 .921 .974 .974 .956 .027
vertebrae T3 .947 .895 .947 .895 .974 .947 .934 .032
vertebrae T4 .949 .897 .923 .923 .949 .949 .932 .021
vertebrae T5 .949 .897 .897 .923 .923 .923 .919 .019
vertebrae T6 .944 .861 .917 .944 .917 .972 .926 .038
vertebrae T7 .872 .846 .897 .897 .897 .897 .885 .021
vertebrae T8 .822 .844 .867 .911 .889 .889 .870 .033
vertebrae T9 .857 .878 .878 .939 .898 .918 .895 .030
Average .914 .880 .887 .924 .901 .902
STD .040 .041 .040 .072 .061 .055
3.2.2 Region-based
LABEL:tab:regionbased-recall-29-regions-reranking and LABEL:tab:regionwise-recall-104-regions-reranking show the retrieval recall for 29 coarse anatomical structures and 104 original TS anatomical structures employing the proposed re-ranking method. Using the re-ranking, the overall performance of all the models has improved. DreamSim performs the best with the average retrieval recall of .$987\pm.027$ and .$987\pm.024$ for 29 and 104 classes, respectively. There are only slight variations between the performance on coarse and all the original TS classes. Similar to the count-based method in the anatomical region retrieval many classes are perfectly retrieved (recall of 1.0). There is a low variation among models and between classes where the highest standard deviation is .064.064.064.064 and .042.042.042.042.
Table 14: Region-based retrieval recall of coarse anatomical regions (29 classes) using HNSW Indexing and re-ranking. In each row, bold numbers represent the best-performing values, while italicized numbers indicate the worst-performing. The separate average and standard deviation (STD) columns are color-coded, with blue indicating the best-performing values and yellow indicating the worst-performing values across different models. Additionally, bold numbers in colored columns represent the best classes in terms of average and standard deviation, while italicized values represent the worst-performing class across the models.
Model DINOv1 DINOv2 DreamSim SwinTrans. ResNet50
Dataset (pre-trained) (ImgNet) (ImgNet) (ImgNet) (RadImg) (Fractaldb) (RadImg) Average STD
adrenal gland 1.000 1.000 1.000 .970 .990 1.000 .993 .012
autochthon .992 1.000 .992 .985 1.000 .992 .994 .006
brain .923 .846 1.000 .923 1.000 1.000 .949 .063
cardiovascular system 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
clavicula .987 .987 1.000 1.000 .974 .987 .989 .010
colon 1.000 1.000 1.000 .981 .981 .981 .991 .010
duodenum .980 .980 1.000 .960 .940 .980 .973 .021
esophagus 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
face .882 .824 .882 .824 .882 .765 .843 .048
femur .965 .977 .988 1.000 1.000 .977 .984 .014
gallbladder .795 .846 .923 .872 .897 .923 .876 .050
gluteus muscles 1.000 1.000 1.000 .988 1.000 .996 .997 .005
hip 1.000 1.000 1.000 1.000 1.000 .989 .998 .005
humerus .954 .954 1.000 .966 .931 .966 .962 .023
iliopsoas 1.000 .990 .990 .981 .981 .990 .989 .007
kidney 1.000 1.000 .990 .990 .980 .990 .992 .007
liver 1.000 1.000 1.000 1.000 1.000 .981 .997 .008
lung 1.000 1.000 1.000 1.000 1.000 .989 .998 .005
pancreas 1.000 1.000 1.000 1.000 .980 1.000 .997 .008
portal and splenic vein .980 .980 .980 .980 .960 .980 .977 .008
rib .999 .999 1.000 1.000 .999 1.000 1.000 .000
sacrum 1.000 1.000 1.000 1.000 1.000 .977 .996 .009
scapula .964 .964 .952 .976 .988 1.000 .974 .018
small bowel .979 1.000 .958 .958 .979 .979 .976 .016
spleen 1.000 1.000 .980 1.000 1.000 1.000 .997 .008
stomach 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
trachea 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
urinary bladder 1.000 1.000 1.000 .977 1.000 .953 .988 .019
vertebrae 1.000 1.000 1.000 .998 .998 .997 .999 .002
Average .979 .977 .987 .977 .981 .979
STD .045 .050 .027 .041 .031 .045
Table 15: Region-based retrieval recall of all TS anatomical regions (104 classes) using HNSW Indexing and re-ranking. In each row, bold numbers represent the best-performing values, while italicized numbers indicate the worst-performing. The separate average and standard deviation (STD) columns are color-coded, with blue indicating the best-performing values and yellow indicating the worst-performing values across different models. Additionally, bold numbers in colored columns represent the best classes in terms of average and standard deviation, while italicized values represent the worst-performing class across the models.
Model DINOv1 DINOv2 DreamSim SwinTrans. ResNet50
Dataset (pre-trained) (ImgNet) (ImgNet) (ImgNet) (RadImg) (Fractaldb) (RadImg) Average STD
adrenal gland left .980 1.000 1.000 .940 1.000 1.000 .987 .024
adrenal gland right 1.000 1.000 1.000 .980 .960 1.000 .990 .017
aorta 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
autochthon left 1.000 1.000 1.000 .985 1.000 .985 .995 .008
autochthon right .985 1.000 .985 .985 1.000 1.000 .992 .008
brain .923 .846 1.000 .923 1.000 1.000 .949 .063
clavicula left .974 .974 1.000 1.000 .949 .974 .979 .019
clavicula right 1.000 1.000 1.000 .974 1.000 1.000 .996 .011
colon 1.000 1.000 1.000 .981 .981 .981 .991 .010
duodenum .980 .980 1.000 .960 .940 .980 .973 .021
esophagus 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
face .882 .824 .882 .824 .882 .765 .843 .048
femur left .956 .956 .978 1.000 1.000 .956 .974 .022
femur right .951 1.000 .951 1.000 1.000 1.000 .984 .025
gallbladder .795 .846 .923 .872 .897 .923 .876 .050
gluteus maximus left 1.000 1.000 1.000 .977 1.000 .977 .992 .012
gluteus maximus right 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
gluteus medius left 1.000 1.000 1.000 .977 1.000 1.000 .996 .009
gluteus medius right 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
gluteus minimus left 1.000 1.000 1.000 .977 1.000 1.000 .996 .009
gluteus minimus right 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
heart atrium left 1.000 1.000 1.000 1.000 1.000 .979 .996 .009
heart atrium right 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
heart myocardium 1.000 1.000 1.000 1.000 1.000 .980 .997 .008
heart ventricle left 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
heart ventricle right 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
hip left 1.000 1.000 1.000 1.000 1.000 .977 .996 .009
hip right 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
humerus left .949 .923 1.000 .974 .897 .897 .940 .042
humerus right .938 .938 1.000 .917 .917 1.000 .951 .039
iliac artery left 1.000 1.000 1.000 .977 1.000 .977 .992 .012
iliac artery right 1.000 1.000 1.000 .977 1.000 .977 .992 .012
iliac vena left 1.000 1.000 1.000 .977 1.000 .977 .992 .012
iliac vena right 1.000 1.000 1.000 .977 1.000 .977 .992 .012
iliopsoas left .981 .981 1.000 .981 .981 .981 .984 .008
iliopsoas right 1.000 .980 .980 .980 .980 1.000 .987 .010
inferior vena cava 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
kidney left .981 .962 .981 .981 .962 .981 .975 .010
kidney right .980 1.000 1.000 .980 1.000 1.000 .993 .011
liver 1.000 1.000 1.000 1.000 1.000 .981 .997 .008
lung lower lobe left 1.000 1.000 1.000 .982 1.000 .965 .991 .015
lung lower lobe right .982 .982 .982 1.000 1.000 .982 .988 .009
lung middle lobe right 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
lung upper lobe left 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
lung upper lobe right .978 .978 .957 1.000 1.000 .978 .982 .016
pancreas 1.000 1.000 1.000 1.000 .980 1.000 .997 .008
portal and splenic vein .980 .980 .980 .980 .960 .980 .977 .008
pulmonary artery .923 .974 .974 1.000 1.000 .974 .974 .028
rib left 1 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
rib left 10 .980 .980 1.000 1.000 .980 1.000 .990 .011
rib left 11 1.000 1.000 1.000 .980 1.000 1.000 .997 .008
rib left 12 .917 .917 1.000 .979 .979 .938 .955 .036
rib left 2 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
rib left 3 1.000 1.000 1.000 1.000 1.000 .976 .996 .010
rib left 4 .975 .975 1.000 1.000 1.000 1.000 .992 .013
rib left 5 1.000 1.000 1.000 .955 1.000 .977 .989 .019
rib left 6 .980 .980 .980 1.000 1.000 1.000 .990 .011
rib left 7 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
rib left 8 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
rib left 9 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
rib right 1 1.000 1.000 1.000 1.000 .947 1.000 .991 .021
rib right 10 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
rib right 11 1.000 1.000 1.000 .980 1.000 1.000 .997 .008
rib right 12 .936 .915 .979 .979 .957 .957 .954 .025
rib right 2 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
rib right 3 .976 1.000 .976 .976 1.000 .951 .980 .018
rib right 4 .975 .975 1.000 1.000 1.000 1.000 .992 .013
rib right 5 1.000 1.000 1.000 .953 1.000 1.000 .992 .019
rib right 6 .980 .980 .980 1.000 1.000 .959 .983 .015
rib right 7 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
rib right 8 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
rib right 9 .980 1.000 .980 1.000 1.000 1.000 .993 .010
sacrum 1.000 1.000 1.000 1.000 1.000 .977 .996 .009
scapula left .976 .976 .951 .976 1.000 1.000 .980 .018
scapula right .953 .953 .953 .977 .977 1.000 .969 .019
small bowel .979 1.000 .958 .958 .979 .979 .976 .016
spleen 1.000 1.000 .980 1.000 1.000 1.000 .997 .008
stomach 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
trachea 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
urinary bladder 1.000 1.000 1.000 .977 1.000 .953 .988 .019
vertebrae C1 .929 .929 .929 .857 .857 .786 .881 .058
vertebrae C2 1.000 .923 1.000 1.000 1.000 1.000 .987 .031
vertebrae C3 .929 .929 1.000 1.000 1.000 1.000 .976 .037
vertebrae C4 .867 .800 .933 .933 .867 .933 .889 .054
vertebrae C5 .895 .842 .895 .842 .842 .842 .860 .027
vertebrae C6 .862 .828 .897 .897 .724 .828 .839 .064
vertebrae C7 1.000 1.000 1.000 1.000 .944 .972 .986 .023
vertebrae L1 1.000 1.000 .980 1.000 .980 1.000 .993 .011
vertebrae L2 .977 .977 .977 .977 1.000 .977 .981 .009
vertebrae L3 .977 1.000 .977 .977 1.000 1.000 .988 .013
vertebrae L4 1.000 .977 .977 1.000 1.000 1.000 .992 .012
vertebrae L5 1.000 .977 .977 1.000 1.000 1.000 .992 .012
vertebrae T1 1.000 1.000 1.000 .973 .919 1.000 .982 .033
vertebrae T10 .980 1.000 1.000 .980 .980 1.000 .990 .011
vertebrae T11 .979 .979 1.000 .979 1.000 1.000 .990 .011
vertebrae T12 1.000 1.000 1.000 .980 1.000 1.000 .997 .008
vertebrae T2 1.000 1.000 1.000 1.000 .974 1.000 .996 .011
vertebrae T3 1.000 .974 1.000 .974 1.000 .974 .987 .014
vertebrae T4 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
vertebrae T5 1.000 .974 .974 1.000 1.000 1.000 .991 .013
vertebrae T6 .917 .972 .972 1.000 .972 .972 .968 .027
vertebrae T7 .949 .949 .949 .974 .974 1.000 .966 .021
vertebrae T8 .955 .977 .955 1.000 1.000 .909 .966 .034
vertebrae T9 .980 .980 .980 .980 1.000 .980 .983 .008
Average .980 .978 .987 .981 .982 .980
STD .036 .042 .024 .033 .042 .041
3.2.3 Localized
Localized Retrieval Recall
LABEL:tab:localized-recall-29-regions-reranking and LABEL:tab:localized-recall-104-regions-reranking show the retrieval recall for 29 coarse anatomical structures and 104 original TS anatomical structures after re-ranking for $L=15$. Re-ranking improved the localization for all the models. DreamsSim is the best-performing model with an average recall of .$955\pm.062$ for coarse anatomical structures and .$956\pm.055$ for original TS classes. Although the retrieval is lower compared to region-based or volume-based evaluation, it is still high which shows that the pretrained vision embeddings not only can retrieve similar cases but also can localize the corresponding region of interest.
Table 16: Localized retrieval recall of coarse anatomical regions (29 classes) using HNSW Indexing and re-ranking, $L=15$. In each row, bold numbers represent the best-performing values, while italicized numbers indicate the worst-performing. The separate average and standard deviation (STD) columns are color-coded, with blue indicating the best-performing values and yellow indicating the worst-performing values across different models. Additionally, bold numbers in colored columns represent the best classes in terms of average and standard deviation, while italicized values represent the worst-performing class across the models.
Model DINOv1 DINOv2 DreamSim SwinTrans. ResNet50
Dataset (pre-trained) (ImgNet) (ImgNet) (ImgNet) (RadImg) (Fractaldb) (RadImg) Average STD
adrenal gland .980 .980 .980 .940 .860 .930 .945 .047
autochthon .992 .985 .985 .985 1.000 .977 .987 .008
brain .923 .846 1.000 .923 .923 1.000 .936 .058
cardiovascular system 1.000 .997 .998 .998 .986 .993 .995 .005
clavicula .961 .987 .987 .961 .909 .961 .961 .028
colon .962 1.000 1.000 .981 .925 .925 .965 .035
duodenum .920 .920 .920 .920 .800 .920 .900 .049
esophagus .946 1.000 1.000 .982 .982 1.000 .985 .021
face .824 .824 .765 .824 .824 .765 .804 .030
femur .930 .965 .965 .977 .942 .977 .959 .019
gallbladder .513 .667 .769 .615 .641 .667 .645 .083
gluteus muscles 1.000 1.000 1.000 .988 .988 .996 .995 .006
hip 1.000 1.000 1.000 1.000 .989 .989 .996 .006
humerus .759 .828 .874 .782 .701 .805 .791 .059
iliopsoas 1.000 .990 .990 .971 .981 .990 .987 .010
kidney .980 .961 .961 .951 .931 .951 .956 .016
liver .944 .981 .963 1.000 .926 .963 .963 .026
lung .977 .992 .996 .973 .966 .973 .980 .012
pancreas .960 .940 .900 .940 .760 .960 .910 .077
portal and splenic vein .920 .960 .920 .940 .760 .940 .907 .073
rib .998 .999 .988 .989 .956 .991 .987 .016
sacrum 1.000 1.000 1.000 .977 1.000 .977 .992 .012
scapula .964 .964 .952 .964 .964 .940 .958 .010
small bowel .979 1.000 .958 .958 .979 .979 .976 .016
spleen 1.000 1.000 .920 .940 .900 1.000 .960 .046
stomach .980 .980 .941 .961 .784 1.000 .941 .079
trachea .976 1.000 1.000 .976 1.000 .951 .984 .020
urinary bladder .977 .930 .977 .953 .953 .953 .957 .018
vertebrae .998 .997 .995 .990 .985 .987 .992 .005
Average .944 .955 .955 .943 .907 .947
STD .099 .075 .062 .079 .099 .076
Table 17: Localized retrieval recall of all TS anatomical regions (104 classes) using HNSW Indexing and re-ranking, $L=15$. In each row, bold numbers represent the best-performing values, while italicized numbers indicate the worst-performing. The separate average and standard deviation (STD) columns are color-coded, with blue indicating the best-performing values and yellow indicating the worst-performing values across different models. Additionally, bold numbers in colored columns represent the best classes in terms of average and standard deviation, while italicized values represent the worst-performing class across the models.
Model DINOv1 DINOv2 DreamSim SwinTrans. ResNet50
Dataset (pre-trained) (ImgNet) (ImgNet) (ImgNet) (RadImg) (Fractaldb) (RadImg) Average STD
adrenal gland left .960 .980 .940 .860 .840 .860 .907 .060
adrenal gland right .940 .940 .940 .900 .820 .940 .913 .048
aorta 1.000 1.000 1.000 1.000 1.000 .984 .997 .007
autochthon left 1.000 .985 1.000 .985 1.000 .954 .987 .018
autochthon right .985 .985 .969 .985 1.000 .985 .985 .010
brain .923 .846 1.000 .923 .923 1.000 .936 .058
clavicula left .949 .974 .974 .949 .846 .949 .940 .048
clavicula right .947 1.000 1.000 .947 .921 .974 .965 .032
colon .962 1.000 1.000 .981 .925 .925 .965 .035
duodenum .920 .920 .920 .920 .800 .920 .900 .049
esophagus .946 1.000 1.000 .982 .982 1.000 .985 .021
face .824 .824 .765 .824 .824 .765 .804 .030
femur left .911 .933 .956 .978 .933 .956 .944 .023
femur right .951 1.000 .927 .976 .951 1.000 .967 .030
gallbladder .513 .667 .769 .615 .641 .667 .645 .083
gluteus maximus left 1.000 1.000 1.000 .977 .977 .977 .988 .013
gluteus maximus right 1.000 1.000 .977 .977 1.000 1.000 .992 .012
gluteus medius left 1.000 1.000 1.000 .977 .977 .977 .989 .012
gluteus medius right 1.000 1.000 1.000 1.000 1.000 1.000 1.000 .000
gluteus minimus left .977 1.000 1.000 .977 .953 .977 .981 .018
gluteus minimus right 1.000 1.000 1.000 1.000 .976 1.000 .996 .010
heart atrium left .936 1.000 1.000 .979 .936 .872 .954 .049
heart atrium right .980 .939 1.000 .959 .918 .918 .952 .033
heart myocardium .980 1.000 1.000 1.000 .918 .959 .976 .033
heart ventricle left .980 .980 1.000 .980 .918 .959 .969 .028
heart ventricle right 1.000 .980 1.000 .980 .959 .959 .980 .018
hip left 1.000 1.000 1.000 1.000 .977 .977 .992 .012
hip right 1.000 1.000 1.000 .977 .977 1.000 .992 .012
humerus left .718 .795 .795 .795 .692 .769 .761 .045
humerus right .688 .792 .833 .688 .667 .792 .743 .071
iliac artery left 1.000 1.000 1.000 .977 .955 .955 .981 .022
iliac artery right 1.000 .977 1.000 .955 .977 .955 .977 .020
iliac vena left 1.000 .977 1.000 .955 .955 .977 .977 .020
iliac vena right 1.000 .977 1.000 .955 .955 .977 .977 .020
iliopsoas left .981 .981 1.000 .962 .981 .981 .981 .012
iliopsoas right 1.000 .980 .980 .980 .980 1.000 .987 .010
inferior vena cava 1.000 .982 1.000 1.000 .947 .982 .985 .021
kidney left .943 .925 .943 .943 .887 .943 .931 .023
kidney right .939 .959 .939 .939 .918 .918 .935 .015
liver .944 .981 .963 1.000 .926 .963 .963 .026
lung lower lobe left .982 1.000 1.000 .965 .965 .930 .974 .027
lung lower lobe right .964 .964 .964 .964 .946 .929 .955 .015
lung middle lobe right .959 .980 .980 .980 .959 .939 .966 .017
lung upper lobe left .946 .982 1.000 .982 .964 .964 .973 .019
lung upper lobe right .935 .957 .935 .935 .935 .957 .942 .011
pancreas .960 .940 .900 .940 .760 .960 .910 .077
portal and splenic vein .920 .960 .920 .940 .760 .940 .907 .073
pulmonary artery .923 .974 .923 .949 .923 .872 .927 .034
rib left 1 1.000 1.000 1.000 .974 .974 .974 .987 .014
rib left 10 .961 .961 .922 .961 .843 1.000 .941 .054
rib left 11 1.000 1.000 .980 .922 .843 .961 .951 .060
rib left 12 .854 .917 .917 .854 .688 .854 .847 .084
rib left 2 .975 .950 1.000 .950 .950 .950 .963 .021
rib left 3 .976 .976 .927 .976 .951 .951 .959 .020
rib left 4 .975 .975 1.000 .975 .975 1.000 .983 .013
rib left 5 .955 .977 .932 .955 .955 .932 .951 .017
rib left 6 .940 .980 .980 .980 .920 .960 .960 .025
rib left 7 .939 1.000 1.000 .980 .939 1.000 .976 .030
rib left 8 .961 1.000 .980 1.000 .902 1.000 .974 .039
rib left 9 .941 .980 .941 .980 .922 .980 .958 .026
rib right 1 1.000 1.000 1.000 .974 .895 .974 .974 .041
rib right 10 1.000 .961 .980 .961 .843 1.000 .958 .059
rib right 11 1.000 1.000 .961 .922 .843 .941 .944 .059
rib right 12 .851 .894 .851 .915 .660 .872 .840 .092
rib right 2 1.000 .974 1.000 .974 .974 .974 .983 .013
rib right 3 .927 1.000 .927 .951 .951 .878 .939 .040
rib right 4 .950 .975 .950 .950 .975 1.000 .967 .020
rib right 5 .953 .953 .977 .953 .930 .977 .957 .018
rib right 6 .959 .980 .980 1.000 .918 .918 .959 .034
rib right 7 .959 1.000 1.000 .980 .939 1.000 .980 .026
rib right 8 .980 .980 .980 1.000 .922 1.000 .977 .029
rib right 9 .941 1.000 .941 .980 .843 .980 .948 .056
sacrum 1.000 1.000 1.000 .977 1.000 .977 .992 .012
scapula left .976 .976 .951 .976 .976 .927 .963 .020
scapula right .953 .953 .953 .907 .953 .953 .946 .019
small bowel .979 1.000 .958 .958 .979 .979 .976 .016
spleen 1.000 1.000 .920 .940 .900 1.000 .960 .046
stomach .980 .980 .941 .961 .784 1.000 .941 .079
trachea .976 1.000 1.000 .976 1.000 .951 .984 .020
urinary bladder .977 .930 .977 .953 .953 .953 .957 .018
vertebrae C1 .929 .929 .929 .857 .786 .786 .869 .070
vertebrae C2 1.000 .923 1.000 1.000 .923 1.000 .974 .040
vertebrae C3 .929 .786 .929 .929 .929 .929 .905 .058
vertebrae C4 .800 .667 .800 .800 .733 .867 .778 .069
vertebrae C5 .789 .684 .895 .789 .684 .684 .754 .086
vertebrae C6 .828 .724 .897 .724 .448 .724 .724 .153
vertebrae C7 .917 .972 1.000 .972 .667 .917 .907 .123
vertebrae L1 .980 .939 .898 .918 .673 .939 .891 .110
vertebrae L2 .932 .886 .955 .886 .841 .841 .890 .046
vertebrae L3 .953 .930 .930 .953 .953 .930 .942 .013
vertebrae L4 .977 .932 .909 1.000 .932 .977 .955 .035
vertebrae L5 .977 .930 .977 .977 .884 .953 .950 .037
vertebrae T1 .946 .946 .946 .946 .649 .919 .892 .120
vertebrae T10 .837 .918 .857 .878 .796 .898 .864 .044
vertebrae T11 .729 .875 .771 .854 .750 .792 .795 .058
vertebrae T12 .920 .940 .920 .900 .740 .880 .883 .073
vertebrae T2 .974 .974 1.000 .974 .842 .974 .956 .057
vertebrae T3 .974 .974 1.000 .921 .947 .947 .961 .028
vertebrae T4 1.000 1.000 1.000 .872 .923 .949 .957 .053
vertebrae T5 1.000 .949 .974 .949 .974 .897 .957 .035
vertebrae T6 .889 .972 .972 .972 .917 .861 .931 .049
vertebrae T7 .923 .949 .923 .897 .923 .872 .915 .026
vertebrae T8 .864 .932 .909 .864 .818 .795 .864 .052
vertebrae T9 .857 .878 .959 .857 .816 .939 .884 .054
Average .944 .951 .956 .940 .890 .935
STD .075 .071 .055 .066 .105 .070
Localization-ratio
The localization-ratio is calculated based on (7). LABEL:tab:localiation-ratio-29-regions-reranking and LABEL:tab:localization-ratio-104-regions-reranking demonstrate the localization-ratio for 29 coarse and 104 original TS anatomical regions after re-ranking for $L=15$. The best-performing embedding is still DreamSim with a localization-ratio of .$837\pm.159$ and .$790\pm.142$ for 29 coarse and 104 original TS classes. After re-ranking the overall localization-ratio is reduced (previously, .$864\pm.0145$ and .$803\pm.130$, respectively).
Table 18: Localization-ratio of coarse anatomical regions (29 classes) using HNSW Indexing and re-ranking, $L=15$. In each row, bold numbers represent the best-performing values, while italicized numbers indicate the worst-performing. The separate average and standard deviation (STD) columns are color-coded, with blue indicating the best-performing values and yellow indicating the worst-performing values across different models. Additionally, bold numbers in colored columns represent the best classes in terms of average and standard deviation, while italicized values represent the worst-performing class across the models.
Model DINOv1 DINOv2 DreamSim SwinTrans. ResNet50
Dataset (pre-trained) (ImgNet) (ImgNet) (ImgNet) (RadImg) (Fractaldb) (RadImg) Average STD
adrenal gland .697 .607 .712 .537 .523 .523 .600 .087
autochthon .970 .956 .951 .959 .926 .939 .950 .016
brain .903 .759 .990 .846 .651 .785 .822 .118
cardiovascular system .945 .977 .989 .968 .983 .962 .971 .016
clavicula .819 .805 .841 .764 .619 .848 .783 .086
colon .874 .886 .952 .808 .891 .843 .875 .049
duodenum .645 .627 .636 .656 .517 .697 .630 .060
esophagus .920 .956 .957 .910 .887 .920 .925 .027
face .682 .702 .569 .729 .561 .576 .637 .076
femur .896 .881 .954 .921 .861 .847 .894 .040
gallbladder .224 .376 .356 .268 .321 .320 .311 .056
gluteus muscles 1.000 .992 1.000 .977 .960 .978 .984 .016
hip .968 .929 .970 .924 .795 .932 .920 .064
humerus .529 .554 .638 .550 .417 .525 .535 .071
iliopsoas .926 .883 .942 .869 .926 .887 .905 .029
kidney .770 .793 .819 .826 .709 .831 .791 .047
liver .804 .788 .742 .737 .604 .770 .741 .072
lung .970 .948 .980 .929 .876 .906 .935 .039
pancreas .772 .721 .773 .720 .505 .651 .690 .101
portal and splenic vein .664 .613 .643 .560 .449 .525 .576 .081
rib 1.000 .982 .992 .970 .918 .992 .976 .030
sacrum .862 .844 .892 .882 .830 .883 .866 .025
scapula .940 .913 .904 .806 .816 .912 .882 .056
small bowel .925 .849 .860 .885 .949 .918 .897 .039
spleen .775 .757 .708 .715 .579 .669 .700 .070
stomach .816 .689 .784 .719 .557 .686 .708 .091
trachea .881 .899 .862 .767 .816 .859 .847 .048
urinary bladder .702 .595 .845 .780 .662 .668 .709 .090
vertebrae 1.000 1.000 1.000 1.000 .981 .970 .992 .013
Average .823 .803 .837 .792 .727 .787
STD .169 .157 .159 .164 .195 .171
Table 19: Localization-ratio of all TS anatomical regions (104 classes) using HNSW Indexing and re-ranking, $L=15$. In each row, bold numbers represent the best-performing values, while italicized numbers indicate the worst-performing. The separate average and standard deviation (STD) columns are color-coded, with blue indicating the best-performing values and yellow indicating the worst-performing values across different models. Additionally, bold numbers in colored columns represent the best classes in terms of average and standard deviation, while italicized values represent the worst-performing class across the models.
Model DINOv1 DINOv2 DreamSim SwinTrans. ResNet50
Dataset (pre-trained) (ImgNet) (ImgNet) (ImgNet) (RadImg) (Fractaldb) (RadImg) Average STD
adrenal gland left .673 .573 .600 .507 .480 .387 .537 .101
adrenal gland right .647 .467 .580 .547 .300 .473 .502 .120
aorta .940 .929 .923 .902 .874 .814 .897 .047
autochthon left .985 .959 .964 .954 .903 .897 .944 .035
autochthon right .933 .964 .928 .954 .908 .908 .932 .023
brain .923 .769 1.000 .897 .667 .821 .846 .119
clavicula left .744 .803 .838 .752 .658 .812 .768 .065
clavicula right .816 .825 .868 .754 .667 .860 .798 .076
colon .881 .855 .931 .818 .899 .855 .873 .039
duodenum .647 .700 .547 .673 .500 .713 .630 .087
esophagus .929 .982 .929 .881 .869 .899 .915 .041
face .608 .706 .549 .706 .549 .647 .627 .071
femur left .881 .896 .956 .911 .837 .859 .890 .041
femur right .951 .951 .927 .951 .919 .886 .931 .026
gallbladder .162 .453 .350 .282 .282 .333 .311 .096
gluteus maximus left .953 .938 .946 .946 .922 .953 .943 .012
gluteus maximus right .946 .930 .930 .946 .938 .961 .942 .012
gluteus medius left .947 .924 .909 .864 .864 .917 .904 .034
gluteus medius right .899 .907 .946 .899 .837 .946 .906 .040
gluteus minimus left .806 .899 .798 .907 .938 .922 .879 .061
gluteus minimus right .865 .841 .802 .905 .881 .976 .878 .060
heart atrium left .645 .745 .837 .596 .667 .525 .669 .110
heart atrium right .721 .776 .898 .741 .646 .762 .757 .082
heart myocardium .748 .769 .939 .782 .687 .830 .793 .085
heart ventricle left .776 .701 .918 .741 .619 .816 .762 .102
heart ventricle right .769 .830 .912 .803 .687 .782 .797 .074
hip left .962 .932 .977 .939 .727 .924 .910 .092
hip right .970 .917 .962 .886 .826 .955 .919 .056
humerus left .530 .607 .615 .590 .368 .513 .537 .093
humerus right .521 .528 .625 .472 .417 .549 .519 .071
iliac artery left .894 .841 .962 .902 .712 .894 .867 .085
iliac artery right .818 .902 .977 .902 .742 .894 .872 .081
iliac vena left .886 .848 .939 .848 .735 .917 .862 .072
iliac vena right .871 .879 .955 .871 .773 .841 .865 .059
iliopsoas left .899 .881 .937 .868 .893 .843 .887 .032
iliopsoas right .889 .895 .895 .895 .967 .889 .905 .031
inferior vena cava .936 .865 .836 .906 .836 .883 .877 .039
kidney left .736 .692 .736 .748 .610 .711 .705 .051
kidney right .646 .735 .782 .748 .633 .762 .718 .063
liver .821 .870 .741 .698 .611 .815 .759 .095
lung lower lobe left .860 .889 .883 .895 .754 .754 .839 .067
lung lower lobe right .875 .839 .857 .893 .720 .720 .817 .077
lung middle lobe right .707 .762 .884 .796 .735 .721 .768 .065
lung upper lobe left .917 .887 .946 .839 .839 .815 .874 .051
lung upper lobe right .855 .899 .841 .775 .826 .754 .825 .053
pancreas .753 .780 .787 .733 .527 .660 .707 .099
portal and splenic vein .733 .713 .673 .573 .480 .473 .608 .116
pulmonary artery .735 .718 .735 .692 .632 .530 .674 .080
rib left 1 .833 .825 .912 .702 .737 .904 .819 .086
rib left 10 .850 .732 .739 .778 .621 .712 .739 .075
rib left 11 .732 .778 .797 .693 .647 .758 .734 .056
rib left 12 .611 .535 .729 .549 .465 .528 .569 .091
rib left 2 .833 .858 .858 .667 .633 .792 .774 .099
rib left 3 .789 .805 .813 .797 .797 .756 .793 .020
rib left 4 .917 .867 .892 .742 .842 .783 .840 .067
rib left 5 .788 .818 .773 .780 .773 .735 .778 .027
rib left 6 .807 .807 .847 .713 .753 .720 .774 .054
rib left 7 .803 .837 .952 .844 .714 .776 .821 .080
rib left 8 .791 .810 .830 .765 .634 .804 .772 .071
rib left 9 .850 .850 .758 .850 .712 .745 .794 .063
rib right 1 .868 .842 .886 .693 .675 .886 .808 .098
rib right 10 .843 .712 .699 .771 .634 .712 .729 .071
rib right 11 .758 .863 .791 .706 .562 .673 .725 .104
rib right 12 .596 .596 .667 .596 .355 .504 .552 .110
rib right 2 .863 .786 .838 .624 .701 .821 .772 .092
rib right 3 .780 .780 .829 .748 .797 .659 .766 .059
rib right 4 .892 .883 .867 .767 .850 .817 .846 .047
rib right 5 .806 .853 .798 .767 .791 .674 .782 .060
rib right 6 .803 .741 .837 .728 .741 .694 .757 .052
rib right 7 .823 .844 .898 .823 .735 .714 .806 .069
rib right 8 .758 .765 .843 .804 .641 .784 .766 .069
rib right 9 .817 .856 .752 .817 .641 .765 .775 .076
sacrum .871 .871 .886 .894 .818 .879 .870 .027
scapula left .927 .894 .886 .846 .821 .902 .879 .039
scapula right .922 .915 .922 .752 .744 .899 .859 .087
small bowel .931 .840 .882 .903 .951 .910 .903 .039
spleen .767 .827 .713 .667 .547 .680 .700 .096
stomach .843 .719 .784 .719 .569 .706 .723 .092
trachea .870 .943 .829 .780 .837 .846 .851 .054
urinary bladder .798 .705 .853 .860 .729 .659 .767 .082
vertebrae C1 .524 .595 .762 .476 .548 .571 .579 .098
vertebrae C2 .744 .667 .846 .692 .513 .718 .697 .109
vertebrae C3 .595 .619 .595 .548 .405 .476 .540 .083
vertebrae C4 .356 .311 .400 .400 .444 .222 .356 .080
vertebrae C5 .544 .491 .632 .456 .351 .333 .468 .114
vertebrae C6 .483 .437 .437 .391 .241 .437 .404 .085
vertebrae C7 .676 .722 .667 .537 .426 .657 .614 .111
vertebrae L1 .728 .558 .687 .680 .442 .612 .618 .105
vertebrae L2 .636 .538 .682 .621 .477 .523 .580 .079
vertebrae L3 .721 .496 .566 .488 .667 .698 .606 .103
vertebrae L4 .636 .379 .568 .795 .644 .576 .600 .136
vertebrae L5 .744 .566 .767 .729 .643 .628 .680 .079
vertebrae T1 .622 .604 .658 .739 .315 .721 .610 .154
vertebrae T10 .571 .524 .619 .599 .483 .490 .548 .057
vertebrae T11 .472 .458 .493 .563 .431 .410 .471 .054
vertebrae T12 .527 .527 .613 .587 .533 .513 .550 .040
vertebrae T2 .632 .623 .509 .728 .675 .737 .651 .084
vertebrae T3 .728 .658 .746 .684 .640 .754 .702 .048
vertebrae T4 .761 .778 .761 .624 .658 .641 .704 .070
vertebrae T5 .692 .709 .786 .692 .547 .581 .668 .088
vertebrae T6 .676 .667 .519 .722 .500 .481 .594 .105
vertebrae T7 .650 .624 .709 .581 .453 .590 .601 .086
vertebrae T8 .576 .545 .750 .462 .439 .348 .520 .139
vertebrae T9 .435 .599 .721 .537 .497 .592 .563 .098
Average .764 .754 .790 .736 .662 .720
STD .148 .150 .142 .145 .171 .166
4 Discussion
4.1 Dataset
As depicted in Figure 6, the labels inside the database and query subset (derived from TS train and test set, respectively) are not balanced. This should resemble a pattern as can be observed in future real-world scenarios of image retrieval. At the same time, this imbalance should be kept in mind when reading and interpreting recall values from the provided result tables.
Figure 6: Distribution of the classes in database (a) and query (b) volumes.
Additionally, it is worth noting that the size and shape of organs can impact the probability of correctly predicting a given label by chance. For example, smaller organs can be less likely to collect "by-chance" true positive predictions compared to larger organs. Similarly, organs with elongated shapes aligned with the slice-wise sampling direction can increase the likelihood of "by-chance" hits. A volume and shape-adjusted representation of recall values does not seem reasonable and thus has not been performed in this work. However, organ volume as shown in Figure 7 and Figure 8 should be considered while interpreting result tables.
Figure 10 and Figure 10 present an overview of mean recall for each of the retrieval methods (all models) versus the mean anatomical region size for 29 and 104 classes, respectively. There is no pattern suggesting any correlation between the size of the anatomical region and the average retrieval recall.
Figure 7: Distribution of the size of the anatomical regions for the 29 regions (a) big anatomical regions with a physical size range that exceeds $10^{6}$ $mm^{3}$, (b) mid-size anatomical regions with a physical size in range in the range (.$1:1)\times 10^{6}( .$ $mm^{3}$ and (c) small anatomical regions with a physical size smaller than .$1\times 10^{6}.$ $mm^{3}$.
Figure 8: Distribution of the size of the anatomical regions for the 104 regions (a) big anatomical regions with a physical size range that exceeds $10^{6}$ $mm^{3}$, (b) mid-size anatomical regions with a physical size in range in the range (.$1:1)\times 10^{6}( .$ $mm^{3}$ and (c) small anatomical regions with a physical size smaller than .$1\times 10^{6}.$ $mm^{3}$.
Figure 9: Overview of average recall vs. mean anatomical region size for 29 anatomical regions for (a) slice-wise, (b) volume-based, (c) volume-based and re-ranking, (d) region-based, (e) region-based and re-ranking, (f) localized, (g) localized and re-ranking retrieval.
Figure 10: Overview of average recall vs. mean anatomical region size for 104 anatomical regions for (a) slice-wise, (b) volume-based, (c) volume-based and re-ranking, (d) region-based, (e) region-based and re-ranking, (f) localized, (g) localized and re-ranking retrieval.
4.2 Re-ranking
For the first time, we could successfully adopt and show the feasibility of ColBERT-inspired re-ranking for an image retrieval task. In theory, this shows that CBIR results can be made subject to context-aware re-ranking. This is very important as it provides a conceptual entry point to use the information of a future retrieval solution in the real world. Concretely, observations such as user behavior on a graphical user interface, and temporal or medical relevance can be "factored in" to adjust the search results. Further research will study the advantages and disadvantages of ColBERT-inspired re-ranking. In future works, further insights into balancing computational costs in the context of latency-accuracy trade-offs will be shared.
4.3 Embeddings
It was shown that embeddings generated from self-supervised models are slightly better for image retrieval tasks than those derived from regular supervised models. This is true for coarse anatomical regions with 29 labels (see LABEL:tab:recall-29-regions) as well as fine-granular anatomical regions with 104 regions (see LABEL:tab:recall-104-regions). This is roughly preserved for all modes of retrieval (i.e. slice-wise, volume-based, region-based, and localized retrieval). More generally, the differences in recall across differently pre-trained models (except pre-trained from fractal image) are very small.Practically, the exact choice of the feature extractor should not be noticeable to a potential user in a downstream application. Further, it can be concluded that pre-training on general natural images (i.e. ImageNet) resulted in slightly more performant embedding vectors than domain-specific images (i.e. RadImageNet). This is unexpected and subject to further research.
Although, the model pre-trained of formula-derived synthetic images of fractals (i.e. Fractaldb) showed the lowest recall accuracy the absolute values are surprisingly high considering that the model learned visual primitives out of rendered fractals. This is very encouraging as the Formular-Driven Supervised Learning (FDSL) can easily be extended to very high number of data points per class and also several virtual classes within one family of formulas (Kataoka et al., 2022). Additionally, the mathematical space of formulas for producing visual primitives is virtually infinite and thus it is the subject of further research whether radiology-specific visual primitives can be created that outperform natural image-based pre-training. Again, FDSL does not require the effort of data collection, curation, and annotation. It can scale to a large number of samples and classes which potentially results in a very smooth and evenly covered latent space.
Embeddings derived from DreamSim architecture showed the highest overall retrieval recall in region-based and localized evaluations. DreamSim is an ensemble architecture that uses multiple ViT embeddings with additional fine-tuning using synthetic images. It is plausible that an ensemble approach outperforms single-architecture embeddings (i.e. DINOv1, DINOv2, SwinTransformer, and ResNet50). Therefore, the usage of DreamSim is currently the preferred method of embedding generation.
Worth discussing is an observation that can be found in all tables presenting recall values. Across all model architectures (column) there are usually a few anatomies or regions (i.e. row) that show lower recall on average (see "Average" column). For example, in LABEL:tab:slicewise-recall-28-regions "gallbladder" showed poor retrieval accuracy, whereas in Table LABEL:tab:volumewise-recall-29-regions "brain" and "face" showed lower recall. The observation of isolated low-recall patterns can be seen across all modes of retrieval and aggregation. The authors of this paper cannot provide an explanation, as to why certain anatomies perform worse in certain retrieval configurations but gain high recall in many other retrieval configurations. This will be subject to future research.
Table 20: Summary of the average retrieval recall and standard deviation between classes for 29 anatomical regions, the boldfaced values highlight the highest recall across feature extractors.
Model DINOv1 DINOv2 DreamSim SwinTrans. ResNet50
Dataset (pre-trained) (ImgNet) (ImgNet) (ImgNet) (RadImg) (Fractaldb) (RadImg)
slice-wise .$855\pm.108$$832\pm.118$$\mathbf{.863\pm.107}$$837\pm.114$$51\pm.52$$813\pm.124$
volume-based .$949\pm.072$$932\pm.064$$936\pm.063$$932\pm.067$$939\pm.078$$\mathbf{.952\pm.043}$
volume-based re-ranked .$\mathbf{.967\pm.040}$$961\pm.045$$967\pm.045$$961\pm.049$$962\pm.086$$960\pm.050$
region-based .$977\pm.047$$976\pm.051$$\mathbf{.979\pm.037}$$973\pm.042$$976\pm.033$$978\pm.039$
region-based re-ranked .$979\pm.045$$977\pm.050$$\mathbf{.987\pm.027}$$977\pm.041$$981\pm.031$$979\pm.045$
localized .$932\pm.105$$\mathbf{.941\pm.077}$$936\pm.089$$922\pm.099$$884\pm.111$$908\pm.099$
localized re-ranked .$944\pm.099$$955\pm.075$$\mathbf{.955\pm.062}$$943\pm.079$$907\pm.099$$947\pm.076$
Table 21: Summary of the average retrieval recall and standard deviation between classes for 104 anatomical regions, the boldfaced values highlight the highest recall across feature extractors.
Model DINOv1 DINOv2 DreamSim SwinTrans. ResNet50
Dataset (pre-trained) (ImgNet) (ImgNet) (ImgNet) (RadImg) (Fractaldb) (RadImg)
slice-wise .$784\pm.137$$750\pm.144$$\mathbf{.797\pm.129}$$765\pm.140$$659\pm.172$$726\pm.154$
volume-based .$\mathbf{.923\pm.077}$$887\pm.071$$892\pm.080$$873\pm.082$$856\pm.054$$908\pm.081$
volume-based re-ranked .$914\pm.040$$880\pm.041$$887\pm.040$$\mathbf{.924\pm.072}$$901\pm.061$$902\pm.055$
region-based .$7.9$037 .$972\pm.050$$\mathbf{.983\pm.032}$$978\pm.032$$973\pm.046$$974\pm.042$
region-based re-ranked .$980\pm.036$$978\pm.042$$\mathbf{.987\pm.024}$$981\pm.033$$982\pm.042$$980\pm.041$
localized .$\mathbf{.929\pm.085}$$917\pm.091$$924\pm.078$$906\pm.092$$840\pm.145$$866\pm.118$
localized re-ranked .$944\pm.075$$951\pm.071$$\mathbf{.956\pm.055}$$940\pm.066$$890\pm.105$$935\pm.070$
Figure 11: An overview of three retrieval methods: (a) Slice-wise, where the retrieval is based on one selected slice e.g., the user zooms to a slice and retrieves the most similar slice (b) volume-based, where the retrieval is based on a complete volume, e.g., the user would like to retrieve similar volumes to the volumes under examination or simply filter the database (c) region-based, where the retrieval is based on the selected organ (or sub-volume), e.g, the user zooms in to a specific region and the most similar volume containing that region is retrieved, (d) localized, where the retrieval is based on the selected organ (or sub-volume) but the region in the retrieved volume is localized to the desired organ (or sub-volume), e.g, the user zooms in to a specific region and the system returns the localized region in the retrieved volume.
4.4 Volume-based, Region-based and Localized Retrieval
Since multiple organs (i.e. labels) are present in each query volume, there are essentially different ways in which image retrieval can be performed. The preferred choice depends on the context of the retrieval task in the real world. If the goal is to find a scan out of a database that is most similar to a complete query scan with the entirety of all present organs (think scan-id to scan-id but visual), then volume-based retrieval is the right choice. In contrast, if the experimenter is interested in a particular organ and its most similar counterpart in the database (and all other organs just happen to be in the same scan due to proximity), then region-based retrieval or localized retrieval is advised. Slice-wise retrieval can find the most similar slice of a volume regardless of the number of other slices. This is not usually a practical choice in real scenarios. Figure 11 visualizes the options.
4.5 Localization-ratio
LABEL:tab:localization-ratio-summary shows the average localization-ratio for with and without re-ranking ($L=15$). There is a drop for both 29 coarse and 104 original TS classes after re-ranking. However, both with and without re-ranking the localization-ratio is high, indicating that most of the slices contributing to the final retrieval of the volumes actually contain the desired anatomical region. However, based on (7) the choice of $L$ can impact this measure. Figure 12 shows localization-ratio for different values of $L$ as shown increasing the $L$ to more that $L=15$ decreases the localization-ratio. This indicates that the first highest values in vector $m_{SIM}$ based on (6) points to the exact slices that contain the desired anatomical structure. To improve the re-ranking localization-ratio, further studies can focus on optimizing the $L$ value based on heuristic, organ size, etc.
Figure 12: Average localization-ratio for different L values for re-ranking evaluation using DreamSim model (best-performing embedding) based on (7).
Table 22: Summary of the average localization-ratio and standard deviation between classes for 29 and 104 anatomical regions for $L=15$, the boldfaced values highlight the highest localization-ratio across feature extractors.
Model DINOv1 DINOv2 DreamSim SwinTrans. ResNet50
Dataset (pre-trained) (ImgNet) (ImgNet) (ImgNet) (RadImg) (Fractaldb) (RadImg)
localization-ratio (29 anatomical regions) $0.842\pm 0.161$ $0.834\pm 0.144$ $\mathbf{0.864\pm 0.145}$ $0.835\pm 0.155$ $0.800\pm 0.168$ $0.815\pm 0.152$
localization-ratio (104 anatomical regions) $0.793\pm 0.140$ $0.763\pm 0.148$ $\mathbf{0.803\pm 0.130}$ $0.773\pm 0.143$ $0.722\pm 0.178$ $0.736\pm 0.162$
localization-ratio re-ranked (29 anatomical regions) $0.823\pm 0.169$ $0.803\pm 0.157$ $\mathbf{0.837\pm 0.159}$ $0.792\pm 0.164$ $0.727\pm 0.195$ $0.787\pm 0.171$
localization-ratio re-ranked (104 anatomical regions) $0.764\pm 0.148$ $0.754\pm 0.150$ $\mathbf{0.790\pm 0.142}$ $0.736\pm 0.145$ $0.662\pm 0.171$ $0.720\pm 0.166$
5 Conclusion
Our study establishes a new benchmark for the retrieval of anatomical structures within 3D medical volumes, utilizing the TotalSegmentator dataset to facilitate targeted queries of volumes or sub-volumes for specific anatomical structures. The results highlight the potential of leveraging pre-trained vision embeddings, originally trained on non-medical images, for medical image retrieval across diverse anatomical regions with a wide size range.
We introduced a re-ranking method based on a late interaction model from text retrieval, i.e. ColBERT Khattab and Zaharia (2020). The proposed ColBERT-inspired method enhances the retrieval recall of all anatomical regions. Future investigations can focus on refining and optimizing the computational efficiency of the proposed re-ranking method.
We evaluated the performance of different embeddings pre-trained supervised and self-supervised on medical and non-medical data. The results indicate that pre-training on general natural images (e.g., ImageNet) yields slightly more effective embedding vectors than domain-specific natural images (e.g., RadImageNet). However, given the marginal difference, the choice of embeddings is unlikely to impact the user experience in downstream tasks significantly.
The retrieval of certain anatomical structures, such as the brain and face, demonstrates low recall across all embedding and retrieval methods. Subsequent research can explore the prevalence of such patterns and potential solutions.
This benchmark sets the stage for future advancements in content-based medical image retrieval, particularly in localizing specific organs or areas within scans.
Acknowledgement
The authors like to thank the Bayer team of the internal ML innovation platform for providing compute infrastructure and technical support.
We thank Timothy Deyer and his RadImageNet team for providing the RadImageNet pre-trained model weights for the SwinTransformer architecture.